### v1

# Story Generation from Visual Inputs: Techniques, Related Tasks, and Challenges

Abstract.
Creating engaging narratives from visual data is crucial for automated digital media consumption, assistive technologies, and interactive entertainment. This survey covers methodologies used in the generation of these narratives, focusing on their principles, strengths, and limitations.
The survey also covers tasks related to automatic story generation, such as image and video captioning, and visual question answering, as well as story generation without visual inputs. These tasks share common challenges with visual story generation and have served as inspiration for the techniques used in the field. We analyze the main datasets and evaluation metrics, providing a critical perspective on their limitations.
Visual Story Generation, Image Captioning, Visual Question Answering, Storytelling
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXjournal: CSURjournalvolume: 00journalnumber: 0article: 000publicationmonth: 0ccs: Computing methodologies Natural language generationccs: Computing methodologies Video summarizationccs: Computing methodologies Scene understanding
1. Introduction
Stories are fundamental to human experience, serving as a bridge between imagination and reality (Hühn et al., 2014). They represent the art of telling tales, whether real or imagined, and present a sequence of events that can captivate, inform, and provoke thought. Storytelling has been an essential part of human culture for millennia, serving as a means of communication, entertainment, education, and moral guidance. The art of storytelling has evolved over time, beginning with the practice of narrating stories through spoken words, advancing to written literature, and, more recently, transitioning to digital media formats. Stories have played a crucial role in shaping societies, passing down knowledge and values from one generation to another. Whether told around a campfire, written in books, or shared through videos and interactive platforms, storytelling continues to captivate and connect people, transcending time and technology (Abbott, 2008).
The field of automated story generation has gained attention, fueled by advances in Natural Language Processing (NLP) (Jurafsky and Martin, 2009) and computer vision (Szeliski, 2010). The goal is to create systems capable of generating novel and engaging stories, similar to those produced by humans (Gervas, 2009). With the increasing availability of images and videos, there is a growing interest in generating stories based on this visual data (Krishna et al., 2017). We refer to this task as Visual Story Generation (VSG).
1.1. Definition and challenges
VSG is the task of automatically creating coherent and engaging stories grounded on visual inputs, such as images or videos. It extends beyond simple image captioning or video description, requiring an understanding of complex relationships, temporal progression, and implicit context within the visual data. The generated narrative must, not only accurately represent the visual content, but also encapsulate the essence of the underlying story, thereby crafting an engaging experience for the audience. VSG serves as a robust indicator of machine comprehension of visual content and poses several challenges, some of which are:
  • Understanding visual content: The recognition of objects and scenes in the visual data, as well as understanding the relationships between them is crucial for generating coherent and meaningful narratives (Xu et al., 2015).
  • Temporal coherence: A story is inherently sequential and maintains a temporal order. Therefore, generated narratives should respect time progression, even if not in a linear manner (e.g., they may involve flashbacks, foreshadowing, or other such mechanisms), and ensure that events are narrated in a logically consistent manner.
  • Contextual understanding: Often, elements of a story can be implicit or inferred from the context. Effective generation approaches must be capable of understanding and incorporating implicit information in the narrative.
  • Engagement and creativity: One of the main goals of a story is to engage the audience. Therefore, the generated narrative should not only be accurate and coherent, but also creative and engaging.
  • Evaluation: Determining the effectiveness of a VSG system is a challenge in itself. Metrics originally developed for tasks such as machine translation and image captioning have been adapted for VSG. However, these metrics are not appropriate to capture the creativity and engagement aspects of a story.
1.2. Background and Motivation
The motivations for VSG ] are numerous. For example, VSG can provide an interactive component to media consumption. Users could potentially influence the direction of the story by choosing different sequences of images or videos, creating a unique, personalized narrative. This could be particularly beneficial in areas like interactive gaming and movies, educational platforms, or even in digital marketing, where personalized narratives can greatly enhance user engagement and satisfaction. VSG can also be used in journalism: it can be used to summarize large amounts of visual data, giving insights into the underlying patterns and relationships, and assisting in content creation.
The growing interest in this area is also motivated by its potential to contribute to the understanding of human cognition and creativity. By developing models that emulate human-like storytelling abilities, researchers can gain insights into the cognitive processes involved in story creation, as well as explore the underlying mechanisms of creativity. These approaches can further lead to novel applications, such as more engaging virtual experiences (Fan et al., 2018; Peng et al., 2018).
These use cases represent only a fraction of the potential applications of VSG. As research in this field progresses, we anticipate that more applications will emerge, spanning an even broader range of domains.
1.3. Scope and structure of the survey
We start by defining the core elements of a story: they are essential for understanding the structure and dynamics of a story and served as foundational components for the development of the early story generation systems. Section 2 presents these elements and discusses their significance in storytelling. It also clarifies the distinctions between story, narrative, and plot, illustrating how a single story can be narrated in multiple ways to yield diverse experiences.
We then explore the main datasets used in story generation, and the various metrics used to assess their performance. We address the limitations that current evaluation metrics face in such an open-ended task and discuss potential directions for future research in this regard. Section 3 addresses these issues.
Since they form the background for VSG approaches, we also explore the methods for textual story generation: from the early rule-based systems to deep learning models. We focus on deep learning methodologies, especially transformer architectures and Large Language Models (LLMs) (Bolaños et al., 2015; Ammanabrolu et al., 2019), due to their state-of-the-art performance. Section 4 presents these approaches and emphasizes their contributions and historical significance.
The connection of VSG with computer vision tasks such as Visual Question Answering (VQA) and image/video captioning is also acknowledged, since they share core challenges, such as visual understanding and narrative generation. Thus, they have contributed with useful techniques to the field. Section 5 explores this convergence between computer vision and NLP.
Concerning VSG proper, we aim to present a balanced perspective on current methodologies, emphasizing opportunities and challenges for future research. The inclusion of a detailed analysis of deep learning methodologies with a broader review of additional and earlier techniques ensures broad coverage of the field. Section 6 presents a broad perspective on the state-of-the-art of VSG.
Finally, Section 7 presents the key findings of this survey and proposes directions for future research.
2. The Elements of Stories
The essence of a story lies in its ability to connect a series of events and experiences, often conveyed by a narrator to an audience. These can range from personal intimate stories to larger epics that span cultures and time.
A story can be told in multiple ways using different narratives. These narratives are more than mere descriptions or dramatic enactments. They involve a process of narration where events are selected, arranged, and em-plotted to give meaning to experiences. This structuring of events is what forms the plot or the story’s synopsis.
The terms story, narrative, and plot are often used interchangeably in everyday language, but they have distinct meanings. According to the Oxford English Dictionary (Simpson, 2002) a story is the straightforward, chronological sequence of events, encompassing characters, settings, and actions. The story is the “what” – what happened, who it happened to, where and when it happened. In contrast, the narrative shapes the form or the “how”, adding context, interpretation, perspective, tone, and style of the storytelling, and nuance to the same set of events. The plot is connected to the narrative, referring to the sequence of events as they are presented to the audience. Thus, a single story can be narrated in multiple ways, each having a different plot and offering a distinct experience to the audience.
Narratives can be broadly categorized as linear or nonlinear (Lau and Chen, 2008), depending on their structure. Each type offers a different approach to unfolding a story and has distinct implications for how the audience experiences and interprets it.
Linear narratives follow a straightforward chronological progression, where events unfold in a sequential and cause-effect manner. Linear narratives often have a single protagonist or a group of characters whose actions and decisions drive the plot forward. This structure is commonly used in traditional storytelling formats. Linear narratives provide a sense of progression and closure, making them easy to follow and understand.
Nonlinear narratives deviate from the traditional linear structure by presenting events out of chronological order. They often involve flashbacks, flash-forwards, or parallel storylines that disrupt the temporal sequence. These narratives can create intrigue, suspense, and complexity by challenging the audience’s expectations and requiring them to piece together the fragmented storyline. The film “Cloud Atlas” (Mitchell, 2008) is an example.
Story elements – character, conflict, theme, setting, plot, and mode – are the essential components that form the foundation of a story and are vital for both crafting and analyzing stories. In the following subsections, we detail each of these elements, discussing their significance in shaping the story.
2.1. Character
Characters are the entities through which the story unfolds (DiBattista, 2011). Their decisions and actions are pivotal in propelling the plot forward. Characters are typically people or other beings, fictional or based on a real-life person, and through their actions and speech, they often convey essential motives and emotions.
The way an audience initially perceives a character significantly influences their ongoing perception and emotional connection, such as empathy or antipathy. This initial impression sets the stage for the audience’s expectations regarding the future behavior of the character. Characters who deviate unexpectedly from their established patterns of behavior can create confusion or surprise, challenging the audience’s understanding and expectations.
Narratives usually center around main characters or protagonists. They are the focal points of the story, often encountering central conflicts, undergoing significant personal growth, or acquiring new knowledge. Many stories also have antagonists, adding complexity and tension to the storyline. In traditional stories, the protagonist often takes the form of a hero, a sympathetic figure battling for morally good causes. Conversely, the antagonist may be portrayed as a villain, acting against these good causes or even perpetrating evil.
Narratives often employ a diverse range of character types, each contributing uniquely to the story’s depth and dynamics. The way characters are developed, presented, and evolve throughout the narrative plays a crucial role in engaging the audience and driving the story’s progression.
2.2. Conflict
Conflict is a fundamental element that generates tension and drives the thoughts and actions of characters (Abbott, 2008). The manner in which conflicts are introduced, developed, and resolved plays a significant role in shaping the trajectory of the narrative and the engagement of the audience with the story. A well-crafted conflict not only propels the narrative forward but also enriches the emotional and psychological depth of the characters and the story itself. The resolution of the conflict, typically towards the end of the narrative, is a crucial aspect. It marks the culmination of tension and provides closure to the story’s central problem.
Broadly, conflict encompasses any form of tension that influences characters, while more narrowly, it may refer to the main challenge faced by the protagonist of the story. This primary conflict often leads to additional struggles for the protagonist, such as feelings of anxiety, insecurity, indecisiveness, or other mental difficulties. These struggles can be categorized as secondary or internal conflicts, adding layers to the journey of the character and depth to the story. Longer stories typically involve multiple conflicts. These can be smaller-scale issues that occur alongside or as part of the main conflict, contributing to the complexity of the story. Commonly used conflict types include: “Character vs. Character” (e.g., opposition); “Character vs. Nature” (e.g., natural forces); “Character vs. Society” (e.g., societal norms); “Character vs. Unavoidable Circumstances” (e.g., fate); and “Character vs. Self” (e.g., dilemmas).
2.3. Theme
Themes encompass the major underlying ideas presented by a story, often abstract and open to the audience’s interpretation (Griffith, 2010). Unlike more concrete story elements, themes invite discussion and contemplation, allowing for a range of interpretations and conclusions. They are not definitive statements but rather exploratory concepts that encourage the audience to engage with and reflect on the story’s deeper meanings. As the narrative unfolds, themes can evolve or reveal new dimensions, offering the audience opportunity to develop their own ideas about the messages or insights the story conveys. The subjective nature of themes means that different audience members may arrive at varying interpretations, which might differ from or even go beyond the creator’s original intent. Thus, understanding and incorporating themes in narrative generation allows for a dynamic, thought-provoking storytelling experience that resonates with and challenges the audience. An example of a theme is the elusive nature of the American Dream, as portrayed by F. Scott Fitzgerald in “The Great Gatsby” (Fitzgerald and Reynolds, 1993).
2.4. Setting
The setting is an element that encompasses the time, place, and context in which the story unfolds (Truby, 2007). It includes the physical and temporal surroundings inhabited by the characters, as well as the social or cultural conventions that influence their actions and decisions. The setting can act dynamically, almost like a character itself, possessing specific traits and undergoing changes that directly affect the plot and contribute to the story’s development. This dynamic nature of the setting allows it to evolve over the course of the story, reflecting and impacting the narrative’s progression. The Hogwarts School of Witchcraft and Wizardry in the “Harry Potter” series (Rowling, 2015) is an example of an evolving setting that influences both plot and character development and impacts the progression of the story.
2.5. Plot
The plot is the sequence of events, the plot points, that propels the story from beginning to middle to end (Dibell, 1999). Each plot point is a moment of change that influences the understanding, decisions, and actions of the characters. This progression is often closely tied to the protagonists encountering or realizing a conflict, and their subsequent efforts to resolve it. This journey creates emotional stakes for both the characters and the audience, generating a sense of suspense. The plot’s movement is integral to the narrative, as it not only advances the storyline but also deepens the audience’s emotional investment in the fate of the characters. A single story can have multiple plots, each representing a different sequence of events or perspective. The plot is usually structured through cause and effect, where actions of the characters and events lead to reactions and consequences, forming a chain of scenes that collectively move the story forward.
As an example, in “The Lord of the Rings” by J.R.R. Tolkien (Tolkien, 2012), the plot unfolds as a series of events beginning with Frodo inheriting the One Ring and learning about its power and dangers, followed by, among other events, the Council of Elrond and the formation of the Fellowship, Gandalf’s fall, the breaking of the Fellowship, and Frodo’s decision to, at first, continue alone. These plot points are influenced by character decisions and actions, which lead to various developments and challenges, enhancing the narrative.
2.6. Mode
The mode is a set of choices and techniques employed by the author to frame the narrative (Pinault, 1992). A single story can be narrated using different modes, each offering a different perspective or experience to the audience. The mode includes the scope of the presented information, the style and type of language used, the medium through which the story is conveyed, and the manner and extent of narrative exposition and commentary.
A key component of the mode is the point of view or perspective from which the story is told. A first-person narrative is an example where a character – often the protagonist – narrates the story using personal pronouns like “I” and “me” directly engaging the audience with their perspective. A third-person narrative is another example that avoids personal pronouns, offering a more detached or impersonal viewpoint. These perspectives are examples of narration, a technique that involves the explicit telling of a story through spoken or written commentary. While narration is frequent in written narratives, it can be optional in other forms, like visual or interactive media.
3. Benchmark Datasets and Evaluation Metrics
This section presents and discusses benchmark datasets used for story generation and evaluation metrics that assess the performance of story generation models.
Although not all approaches are data-driven, some of them, especially deep learning models and other machine learning approaches, rely heavily on data to guide the learning process. Various story datasets have been employed in the development of story generation systems by the community.
Table 1 provides a compilation of the datasets documented in the literature for both VSG and textual story generation. Among these datasets, VIST (Huang et al., 2016) is prominently utilized by the research community. This dataset serves as a primary benchmark, and most publications in this field report their experimental results on it. The Visual Writing Prompts dataset (Hong et al., 2023) is noteworthy for its focus on stories that are visually grounded on the characters and objects present in the images. Similarly, the Writing Prompts dataset (Fan et al., 2018) and STORIUM (Akoury et al., 2020) are important because they contain long stories derived from the imagination of the users of the platforms where they were collected.
These datasets allow for the exploration of more complex and imaginative narrative structures, serving as a resource to train models to generate fictional stories.
Dataset No. of Stories Type Notes
NY City (Park and Kim, 2015) 11,863 visual 11,863 blog posts / 78,467 images
Disneyland (Park and Kim, 2015) 7,717 visual 5 7,717 blog posts / 60,545 images
ROCStories (Mostafazadeh et al., 2016) 100,000 visual 5 sentences/story.
LSMDC (Rohrbach et al., 2016) 118,114 visual 202 movies.
VIST (Huang et al., 2016) 50,000 visual 210,819 unique photos
Visual Writing Prompts (Hong et al., 2023) 12,000 visual 5 to 10 images/story, grounded on the characters.
Pororo-SV (Kim et al., 2017) 27,328 visual
YouCook2 (Das et al., 2013) 2,000 visual Average length of 5.26 mins.
Writing Prompts (Fan et al., 2018) 300,000 text 28.4 words/prompt, 734.5 words/story.
CMU movie summary corpus 42,306 visual Includes box office revenue, genre, release date, runtime, character names and actor names.
Wikipedia’s movie plots 34,886 text Summary of the movie plot with metadata.
Sci-fi TV plot summaries (Ammanabrolu et al., 2019) 2,276 text 80+ episodes of TV Series.
STORIUM (Akoury et al., 2020) 5,743 text 19,278 words/story. Annotations for scene entries and cards played to guide the story.
Table 1. Story Generation Datasets.
The evaluation metrics, used to assess the performance of VSG models, provide quantitative measures to gauge the quality and coherence of generated narratives. Most of these metrics were originally developed for machine translation or summarization but have been adopted by the community for other tasks such as image captioning and story generation: instead of comparing a generated text with a reference translation, the generated text is compared with a reference caption in the case of image captioning, or with a reference story in the case of story generation. In the following subsections, we discuss commonly used metrics, including their strengths and limitations.
3.1. Perplexity
Perplexity (Jurafsky and Martin, 2023) evaluates the performance of autoregressive language models. It measures the model’s ability to predict a sequence and is defined as the exponential average negative log-likelihood of a sequence: given a tokenized sequence $X=(x_{1},x_{2},\ldots,x_{N})$, where $N$ is the length of the sequence, the perplexity of $X$ is given by Eq. 1. $P(x_{i}|x_{1},\ldots,x_{i-1})$ denotes the probability of token $x_{i}$ given the preceding token sequence $x_{1},\ldots,x_{i-1}$ as predicted by the model. Perplexity can be interpreted as the geometric mean of the inverse probability of the word sequence, indicating how well the model predicts each subsequent token in the sequence. Lower values indicate better predictive performance, meaning that the model is more effective in capturing the underlying structure of the language.
(1) $\text{Perplexity}(X)=\sqrt[N]{\prod_{i=1}^{N}\frac{1}{P(x_{i}|x_{1},\ldots,x_{i-1})}},$
3.2. BLEU
Bilingual Evaluation Understudy (BLEU) (Papineni et al., 2002) evaluates the quality of generated text by comparing it with one or more reference texts. It was developed for machine translation tasks, but has been adopted for various natural language generation tasks including VSG. BLEU outputs a score between 00 and $1$, indicating the similarity between the generated and reference texts. Higher scores reflect better alignment with the reference: a score of $1$ signifies perfect agreement.
BLEU compares n-grams in the candidate text with the n-grams in the reference texts, computing the precision of these matches, which are independent of their position in the text. The general idea is that the greater the number of matching n-grams, the higher the quality of the text. To mitigate the issue of favoring shorter texts (since they can potentially have higher precision), BLEU incorporates a brevity penalty. This penalty balances precision and recall, ensuring that the metric does not disproportionately favor shorter stories.
In Eq. 2, BP is the brevity penalty (Eq. 3), $w_{n}$ is the weight for each n-gram, $p_{n}$ is the n-gram precision, and $N$ is typically up to four, accounting for the use of unigrams, bigrams, trigrams, and four-grams.
(2) BLEU $=\text{BP}\cdot\exp\left(\sum_{n=1}^{N}w_{n}\log p_{n}\right)$
(3) BP =$=\begin{cases}1&\text{if }c>r\\ e^{(1-r/c)}&\text{if }c\leq r\end{cases}$
While widely used, BLEU has limitations in evaluating stories involving word variations, synonyms, and creative text. Its reliance on exact n-gram matches fails to capture semantic equivalence when different wording or phrasing is used, potentially penalizing stories that are semantically accurate but stylistically different. This results in reduced effectiveness in assessing linguistic richness, idiomatic expressions, and the subtleties of creative language use.
3.3. Self-BLEU
Self-BLEU (Zhu et al., 2018) evaluates diversity within a set of generated data. While the original BLEU was designed to assess the similarity between two texts, Self-BLEU adapts this concept to measure how much each sentence within a generated text resembles the rest of the text. An individual sentence is treated as the hypothesis and all other companion sentences are considered as the reference set. The BLEU score is calculated for each sentence against the aggregated reference set, and the average of these scores constitutes the Self-BLEU score for the entire document. In Eq. 4, $N$ is the number of sentences, $s_{i}$ is the $i$-th sentence, and $S\setminus\{s_{i}\}$ denotes the set of all sentences excluding $s_{i}$.
[…]
A higher Self-BLEU score indicates a lower diversity among the generated sentences, suggesting a potential mode collapse issue within the generative model. Mode collapse occurs when the model produces a narrow range of outputs, repeatedly generating similar types of data while failing to capture the full diversity of the target distribution. Conversely, a lower Self-BLEU score suggests greater diversity and a healthier, more varied generation process. This is particularly relevant for story generation models where diversity in narrative output is crucial for creating engaging and unique stories. High Self-BLEU scores might signal a lack of creativity and variability in the stories produced, whereas lower scores indicate a model’s effectiveness in generating rich, varied narratives.
3.4. METEOR
Metric for Evaluation of Translation with Explicit ORdering (METEOR) (Banerjee and Lavie, 2005) was designed to evaluate machine translation outputs, addressing some limitations of BLEU. It focuses on sentence-level evaluation, aiming to achieve higher correlation with human judgments. The metric relies on the harmonic mean of unigram precision and recall, with emphasis placed on recall to better align with human evaluative tendencies.
The algorithm initiates by aligning the candidate text with the reference text, focusing on unigram mappings. These mappings are constrained such that each unigram in the candidate text aligns with zero or one unigram in the reference text. The alignment process prioritizes configurations with the fewest crossing lines to ensure clarity and coherence.
Precision $P={m}/{w_{t}}$ and recall $R={m}/{w_{r}}$ are the ratios between $m$ (the number of matching unigrams in the candidate text) and either $w_{t}$ (the number of unigrams in the candidate text) or $w_{r}$ (the number of unigrams in the reference text). These metrics are integrated using the harmonic mean, emphasizing recall $F_{mean}=10\cdot P\cdot R/(R+9\cdot P)$. To account for the arrangement of words, METEOR introduces a penalty $p=0.5(c/u_{m})^{3}$, where $c$ is the number of chunks, and $u_{m}$ is the number of mapped unigrams. The final METEOR score for a candidate text is given by $M=(1-p)F_{mean}$.
3.5. CIDEr
Consensus-based Image Description Evaluation (CIDEr) (Vedantam et al., 2014) was developed for image captioning evaluation. It evaluates how well a candidate sentence $c_{i}$ aligns with a consensus among a set of reference sentences $S_{i}=\{s_{i1},\ldots,s_{im}\}$ for an image $I_{i}$. CIDEr involves stemming words in the sentences to their root forms. Each sentence is represented by its n-grams, where an n-gram $\omega_{k}$ is an ordered set of words, and n-grams from one to four words are considered.
CIDEr evaluates the consensus by measuring how frequently n-grams in the candidate sentence appear in the reference sentences. It also assigns lower weights to n-grams common across all images in the dataset, as they are likely less informative. The Term Frequency-Inverse Document Frequency (TF-IDF) (Leskovec et al., 2014) method is employed to weight each n-gram $\omega_{k}$ based on its informativeness, as shown in Eq. 5, where $h_{k}(s_{ij})$ is the frequency of $\omega_{k}$ in sentence $s_{ij}$, and $\Omega$ is the set of all n-grams and $|I|$ is the set of all images in the dataset.
(5) $g_{k}(s_{ij})=\frac{h_{k}(s_{ij})}{\sum_{\omega_{l}\in\Omega}h_{l}(s_{ij})}\log\left(\frac{|I|}{\sum_{I_{p}\in I}\min(1,\sum_{q}h_{k}(s_{pq}))}\right),$
The CIDEr score for n-grams of length $n$ is calculated using the average cosine similarity between the TF-IDF vectors of the candidate and reference sentences, capturing both precision and recall as shown in Eq. 6, where $g_{n}(c_{i})$ and $g_{n}(s_{ij})$ are the TF-IDF vectors for n-grams of length $n$ in the candidate and reference sentences, respectively. $m$ is the number of reference sentences for the candidate image $i$.
(6) $CIDEr_{n}(c_{i},S_{i})=\frac{1}{m}\sum_{j}\frac{g_{n}(c_{i})\cdot g_{n}(s_{ij})}{\|g_{n}(c_{i})\|\|g_{n}(s_{ij})\|},$
To derive the final score, CIDEr combines scores across different n-gram lengths, as shown in Eq. 7, where uniform weights $w_{n}=1/N$ have empirically been found to perform best, and $N$ is typically set to 4.
(7) $CIDEr(c_{i},S_{i})=\sum_{n=1}^{N}w_{n}CIDEr_{n}(c_{i},S_{i}),$
3.6. ROUGE
Recall-Oriented Understudy for Gisting Evaluation (ROUGE) (Lin, 2004) is a recall-oriented measure of the overlap of n-grams between a candidate text and reference texts. It was developed for summarization tasks, but has been used in other natural language generation tasks. ROUGE-N, one of the most common variations of the metric, is the ratio of the number of n-grams co-occurring in the candidate text and the reference texts to the total number of n-grams in the reference texts, as shown in Eq. 8, where $\text{gram}_{n}$ is the n-gram being considered, and $\text{Count}_{\text{match}}(\text{gram}_{n}(S))$ is the maximum number of n-grams co-occurring in a candidate and a set of reference texts.
[…]
As the number of reference texts increases, the denominator in the ROUGE-N formula also increases, which is intuitive since adding more references expands the space of acceptable texts, enhancing the robustness of the evaluation.
For multiple references, ROUGE-N is calculated as the maximum of the pairwise ROUGE-N scores between the candidate text and each reference text (Eq. 9, where $r_{i}$ is a reference text, and $s$ is the candidate text).
[…]
This approach is also extended to other ROUGE metrics such as ROUGE-L (based on the longest common subsequence), ROUGE-W (considers word sequences with gaps), and ROUGE-S (based on co-occurrence statistics of skip-bigrams, i.e., pairs of words in sentence order, allowing arbitrary gaps between them).
3.7. BERTScore
BERTScore (Zhang et al., 2019) is a metric that uses contextual embeddings to evaluate the similarity between a reference sentence $x=\{x_{1},\ldots,x_{k}\}$ and a candidate sentence $\hat{x}=\{\hat{x}_{1},\ldots,\hat{x}_{l}\}$. In story generation, $x$ and $\hat{x}$ are the reference and candidate stories, respectively. Each token in $x$ and $\hat{x}$ is a contextual embedding obtained from models trained on language modeling objectives, enabling the generation of distinct vectors for the same word in different contexts. The similarity between tokens is computed using cosine similarity. For pre-normalized vectors, this calculation simplifies to the inner product of the embeddings, that is, $\text{similarity}(x_{i},\hat{x}_{j})=x_{i}^{\top}\hat{x}_{j}$.
BERTScore computes recall $R_{\text{BERT}}$, precision $P_{\text{BERT}}$, and F1 score $F_{\text{BERT}}$ by matching each token in one sentence to the most similar token in the other sentence. These scores are defined as shown in Eqs. 1011, and 12, respectively.
(10) $R_{\text{BERT}}=\frac{1}{|x|}\sum_{x_{i}\in x}\max_{\hat{x}_{j}\in\hat{x}}x_{i}^{\top}\hat{x}_{j}$
(11) $P_{\text{BERT}}=\frac{1}{|\hat{x}|}\sum_{\hat{x}_{j}\in\hat{x}}\max_{x_{i}\in x}x_{i}^{\top}\hat{x}_{j}$
(12) $F_{\text{BERT}}=\frac{2\cdot P_{\text{BERT}}\cdot R_{\text{BERT}}}{P_{\text{BERT}}+R_{\text{BERT}}}$
BERTScore can incorporate importance weighting using Inverse Document Frequency (IDF) scores, emphasizing rare words which are often more indicative of sentence similarity. Eq. 13 shows the recall with IDF weighting.
(13) $R_{\text{BERT}}^{\text{idf}}=\frac{\sum_{x_{i}\in x}\text{idf}(x_{i})\max_{\hat{x}_{j}\in\hat{x}}x_{i}^{\top}\hat{x}_{j}}{\sum_{x_{i}\in x}\text{idf}(x_{i})}$
To make BERTScore more interpretable, it is rescaled against an empirical lower bound $b$, derived from averaging BERTScores computed on randomly paired sentences from a large corpus. Eq. 14 shows the rescaled BERTScore $\hat{R}_{\text{BERT}}$.
(14) $\hat{R}_{\text{BERT}}=\frac{R_{\text{BERT}}-b}{1-b}$
3.8. Discussion
These metrics guide the development and refinement of visual story generation models by offering insights into their linguistic fluency, relevance to visual content, and overall narrative coherence. They focus on specific linguistic and content-related aspects, providing quantitative measures for precision, recall, and consensus. They are, however, limited by their focus on surface-level textual similarities rather than deep narrative alignment: since a story can be retold in various ways, employing different perspectives while preserving the original message, the corresponding generated text might deviate significantly from the reference text in terms of specific wording.
Huang et al. (Huang et al., 2016) conducted an analysis to evaluate how well BLEU and METEOR correlate with human judgment in the context of VSG according to different metrics. The results in Table 2 show that METEOR, which places a strong emphasis on paraphrasing, has higher correlation with human judgment than BLEU. Nevertheless, in spite of METEOR showing the highest correlation with human judgment, this value is still low, suggesting that a large gap still exists between automatic metrics and human judgment.
Metric METEOR BLEU
Pearson 0.22 (2.8e-28) 0.08 (1.0e-06)
Spearman 0.20 (3.0e-31) 0.08 (8.9e-06)
Kendall 0.14 (1.0e-33) 0.06 (8.7e-08)
Table 2. Correlations of automatic scores against human judgements, with p-values in parentheses (Huang et al., 2016).
To evaluate VSG models, additional metrics that capture the creative aspects of storytelling may be necessary. These metrics might assess the uniqueness, novelty, and diversity of generated narratives, ensuring that models produce stories that go beyond mere data memorization. Coherence metrics could delve deeper into the logical flow of events, ensuring that generated narratives maintain a consistent and plausible storyline.
The proliferation of diverse LLMs, such as the GPT (OpenAI, 2024c, b, a), Llama (Touvron et al., 2023; Meta, 2024), Mistral (team, 2023), and Claude (Anthropic, 2024b, a, 2024) families, has led to the development of systematic scoring systems, aiming for objective assessments across a wide array of language understanding tasks. These scoring systems include Multi-turn Benchmark (MT-Bench) (Zheng et al., 2023) and Chatbot Arena (Zheng et al., 2023) briefly described below. While they were initially developed for the evaluation of chatbots by human evaluators, they have been adapted to operate using a LLM as the evaluator achieving results close to those obtained by human evaluators.
As LLMs continue to improve, they show potential in replacing human annotators in many tasks (Gilardi et al., 2023). Metrics to evaluate stories could be developed based on LLMs (Zheng et al., 2023), possibly following the underlying principles discussed above.
MT-Bench is a benchmark tailored to test the multi-turn conversation and instruction-following capabilities of LLM-based chat assistants. It includes 80 high-quality multi-turn questions distributed across diverse categories such as writing, roleplay, reasoning, and knowledge domains. In the MT-Bench evaluation process, the user is presented with two distinct conversations, each generated by a different LLM-based chat assistant and is then tasked with deciding which assistant, A, B, or indicating a tie, better followed the instructions and answered the questions.
Chatbot Arena introduces a crowd-sourcing approach, fostering anonymous battles between LLM-based chatbot models. Users interact with two anonymous models simultaneously, posing identical questions to both and subsequently voting for the model that provides the preferred response. The Elo rating system (Zheng et al., 2023) is employed to rank the models based on their performance in these battles. This rating system assigns a numerical score to each model based on their battle outcomes, adjusting these scores higher or lower after each encounter to reflect their relative ability to satisfy user queries effectively compared to their opponents.
In summary, while current metrics provide insights into specific dimensions of VSG, the development of new metrics may be necessary to holistically evaluate creativity, coherence, and emotional engagement, ensuring a more comprehensive understanding of the capabilities and limitations of these models. We anticipate that LLMs can play a role in the development of such metrics.
4. Story Generation
The use of Artificial Intelligence (AI) for automated story generation has been explored since the 1960s and 1970s. Early approaches include rule-based systems (Section 4.1), plan-based methods (Section 4.3), grammar-based systems (Section 4.2), and case-based reasoning systems (Section 4.4). In recent times, with the availability of large datasets and computational resources the these methods have been replaced by deep learning models: Recurrent Neural Networks (RNNs) (Section 4.5) which have been widely explored for story generation tasks; and LLMs (Section 4.6) which currently achieve state-of-the-art performance in various natural language generation tasks.
4.1. Rule-based Methods
Rule-based methods in narrative generation use a predefined set of rules to create stories. These rules govern the relationships between narrative elements, ensuring consistency within the narrative, aligning with the desired outcomes and narrative styles or genres. However, it is essential to acknowledge the potential limitations in flexibility and creativity, as the story’s possibilities are bounded by the predefined rules and the system’s ability to simulate complex interpersonal dynamics and character development.
TALE-SPIN (1977) (Meehan, 1977) was an influential rule-based method that initiated its storytelling process by presenting a central problem to the main character, derived from foundational states. As the narrative unfolds, the character may encounter various sub-problems and unintended side effects, which also require resolution. The complexity of these challenges can be adjusted by the user, influencing the story’s length and intricacy. This capability to modulate narrative complexity allows for a range of storytelling from straightforward tales to more complex narratives, depending on the inter-character dynamics established by the user. A pivotal aspect of narrative construction of TALE-SPIN is its nuanced handling of character relationships which are essential for the progression and richness of the story. These relationships are categorized into various dimensions such as competition, dominance, familiarity, affection, trust, deceit, and indebtedness. These relationships not only set preconditions for the actions of the characters but also emerge as outcomes of their interactions within the narrative. For instance, the act of one character giving a gift to another can lead to the recipient feeling a sense of affection towards the giver, thereby influencing subsequent narrative developments. Fig. 1 shows an example of a story generated by TALE-SPIN.
[…]
Figure 1. An example story by TALE-SPIN (Meehan, 1977).
IDtension (2003) (Szilas, 2003) merges dramatic storytelling with interactive fiction. This architecture is designed to balance narrative consistency with the creative spontaneity essential for immersive interactive experiences. IDtension is composed of five foundational modules: the world of the story, narrative logic, narrative sequencer, user model, and the theater. Each module plays a role in crafting a narrative that responds and evolves based on user interactions, maintaining an equilibrium between narrative depth and interactive freedom. The world of the story module serves as the narrative’s backbone, housing the essential entities and their states, thereby setting the stage for the unfolding drama. The narrative logic then calculates all possible character actions derived from the narrative ecosystem. The narrative sequencer filters through actions to highlight those with the greatest narrative value, ensuring that each user decision is met with consequences that propel the story forward. The user model captures the user’s current narrative context, guiding the sequencer in tailoring the narrative flow to enhance user engagement. The theater is responsible for displaying the action(s), and manages the interaction between the computer and the user. Fig. 2 shows an example of a story generated by IDtension.
[…]
Figure 2. An example story by IDtension (Szilas, 2003) showing the telegraphic nature of the text.
4.2. Grammar-based Methods
Grammar-based methods in narrative generation, a specialized subset of rule-based approaches, employ formal grammars, such as context-free grammars, to systematically construct stories. The vocabulary elements can be narrative elements like the ones described in the previous section. In contrast with rule-based methods, which emphasize logical consistency and specific narrative outcomes through predefined rules guiding character and plot interactions, these grammars consist of a set of production rules that define how various elements of a story can be combined to form coherent narratives. This method emphasizes linguistic correctness and structural coherence, making it suitable for applications where the narrative needs to adhere to specific stylistic or syntactic standards. However, the creativity and variety of the narratives generated are often limited by the constraints of the grammar used. Grammar-based systems are particularly effective in scenarios where the narrative needs to follow a well-defined and consistent structure.
Different grammars have been proposed that capture narrative patterns such as the three-act structure (Aristotle et al., 1997), Propp’s morphology (Propp et al., 2010), Freytag’s pyramid (Freytag and MacEwan, 1894), the hero’s journey monomyth (Campbell, 2008), among others. Grammars can be designed for specific genres or domains.
Examples of grammar-based methods include systems developed by Grimes (Ryan, 2017) and Gervás (Gervás, 2013). Grimes’ system was one of the pioneering approaches in the 1960s to harness computational methods for automatic story generation. The system would randomly select elements from Propp’s list (Propp, 1968) and then order them properly to generate a coherent story. This was a significant achievement, considering the computational limitations of the time. This system employed a grammar-based technique, drawing inspiration from Propp’s morphology, to craft structured narratives. Gervás (2013), decades after Grimes’ system, computationally modeled Russian folktales. His approach involved representing Propp’s morphology by analyzing character roles and narrative functions. This was implemented as a generative grammar that arranges plot elements according to a chronological sequence of narrative functions.
4.3. Plan-based Methods
Plan-based methods for narrative generation involve using planning algorithms to create a storyline. In this approach, the system formulates a narrative goal and then plans a sequence of events or actions that characters can take to achieve that goal. This method allows for dynamic storytelling, as the system can adapt to different narrative goals and constraints. However, incoherent stories can be generated if the planning algorithm is not sophisticated enough. Plan-based methods are useful for creating stories that require logical progression and can adapt to changing story requirements. Examples of plan-based methods include MEXICA (Pérez y Pérez and Sharples, 2001) and Mimesis (Young et al., 2004).
4.3.1. MEXICA
MEXICA (1999) (Pérez y Pérez and Sharples, 2001) used creative storytelling to generate Aztec mythology tales. Character-based planning drove narrative progression based on changes in character emotions extracted from a case library.
MEXICA operates through two main processes. Initially, it builds knowledge from a collection of predefined stories, using this information to create schemas in its Long-Term Memory (LTM) that encapsulate content and rhetorical knowledge specific to a genre or form of writing. Users can add or modify stories in the system’s archive using a specialized language designed for defining these stories (PSDL) and the actions within them (SADL).
The MEXICA storytelling core is an engagement-reflection cycle. In the engagement phase, MEXICA generates stories by planning a sequence of actions, each with associated post-conditions that alter the story world (e.g., emotional links between characters or changes in their physical positions). These post-conditions are not explicitly stated in the story but guide the narrative’s direction. For instance, the action “princess cures jaguar knight’s injuries” not only progresses the plot but also establishes an emotional link (gratitude) between the characters, influencing future actions.
Characters in MEXICA possess individual “story-world contexts”, which are dynamic records of the post-conditions affecting them. These contexts ensure that each character’s storyline reflects their unique experiences and interactions in the narrative, allowing for complex and divergent storylines based on perspectives and relationships of the characters.
When MEXICA reaches a narrative impasse, that is, when it cannot proceed based on the current story-world context, it enters the reflection phase. Here, it analyzes previous stories for inspiration to overcome the block, assessing novelty and interestingness by comparing the current story’s content and dramatic tension with those of past narratives. This process allows MEXICA to break impasses and continue with story generation, thereby simulating a human writer’s strategy of seeking external inspiration to resolve creative obstacles. Fig. 3 shows a story generated by MEXICA. This example shows how limited pronoun variation and repetitive use of character names cause the narrative to be less fluid.
[…]
Figure 3. An example story by MEXICA (Pérez y Pérez and Sharples, 2001).
4.3.2. Mimesis
Mimesis (2004) (Young et al., 2004) addresses a fundamental challenge in game development: creating compelling and context-sensitive behaviors for characters and objects. Its architecture employs a structured approach to generate and execute narrative plans within interactive game environments. The process begins when the world planner receives a request from the game engine, detailing a narrative problem that requires resolution, such as achieving specified narrative objectives using a predetermined set of actions. In response, the world planner formulates a plan, a structured dataset that sequences actions for characters and other game entities to fulfill the narrative goals outlined in the request. Following the creation of the story world plan, the discourse planner manages non-character elements like camera movements and background music, enhancing the narrative unfolding within the game. The discourse planner then uses the world plan to craft a narrative plan that encapsulates all system and user activities planned in response to the game engine’s initial request. The final step involves the execution manager, which receives the narrative plan and builds a Directed Acyclic Graph (DAG) to map temporal dependencies among the actions. It then navigates through the DAG, initiating ready actions, commanding the game engine to execute them, and monitoring their outcomes to ensure successful completion or to record any failures. We do not include an example story of Mimesis as it is a system designed to generate stories in interactive game environments.
4.4. Case-based Reasoning Systems
Case-based reasoning approaches generate new stories by retrieving and adapting prior story plots stored in a memory of narrative cases. When a new story situation is defined, including elements like setting, characters, and author goals, this situation is matched against the narrative case library to find the most similar prior plots. One or more retrieved plots are adapted to the new situation by modifying characters, events and descriptions, creating a new plot that becomes the framework for the new story (it can also be added to the library for future reuse). This approach allows generating novel narratives by reusing and building upon successful past story structures.
Case-based narrative systems must define how to represent, index, match, adapt, and learn from story cases. Large case libraries, similarity metrics, and adaptation rules enable generating a wide range of creative plot variations.
Examples of case-based reasoning systems include MINSTREL (1993) (Turner, 1993) and Story Plot Generation (2004) (Gervás et al., 2004). The latter uses existing stories and case-based reasoning to create new stories based on user queries. MINSTREL is a creative storytelling system that generated original stories about King Arthur. It combines case-based goal planning with models of drama and emotional responses. Fig. 4 shows an example of a story created by the system.
[…]
It was the spring of 1089, and a knight named Lancelot returned to Camelot from elsewhere. Lancelot was hot tempered. Once, Lancelot lost a joust. Because he was hot tempered, Lancelot wanted to destroy his sword. Lancelot struck his sword. His sword was destroyed.
One day, a lady of the court named Andrea wanted to have some berries. Andrea went to the woods. Andrea had some berries because Andrea picked some berries. Lancelot’s horse moved Lancelot to the woods. This unexpectedly caused him to be near Andrea. Because Lancelot was near Andrea, Lancelot saw Andrea. Lancelot loved Andrea.
Some time later, Lancelot’s horse moved Lancelot to the woods unintentionally, again causing him to be near Andrea. Lancelot knew that Andrea kissed with a knight named Frederick because Lancelot saw that Andrea kissed with Frederick. Lancelot believed that Andrea loved Frederick. Lancelot loved Andrea. Because Lancelot loved Andrea, Lancelot wanted to be the love of Andrea. But he could not because Andrea loved Frederick. Lancelot hated Frederick. Because Lancelot was hot tempered, Lancelot wanted to kill Frederick. Lancelot went to Frederick. Lancelot fought with Frederick. Frederick was dead.
Andrea went to Frederick. Andrea told Lancelot that Andrea was siblings with Frederick. Lancelot believed that Andrea was siblings with Frederick. Lancelot wanted to take back that he wanted to kill Frederick, but he could not because Frederick was dead. Lancelot hated himself. Lancelot became a hermit. Frederick was buried in the woods. Andrea became a nun.
[…]
Figure 4. An example story by MINSTREL (Turner, 1993).
MINSTREL narratives, while coherent, lack several elements that would make them more engaging. The story’s phrasing is repetitive, which, along with the lack of descriptive detail, hampers the reader’s ability to fully immerse themselves in the world or connect with the characters. It also misses opportunities to demonstrate the emotions and motivations of the characters through their actions or dialogues, rather than through plain statements.
4.5. Recurrent Neural Networks
Recurrent Neural Network (RNN) are a deep learning models for sequential data processing, with applications ranging from language modeling to image captioning and text generation, that influenced later models. Sequential processing in RNNs involves handling sentences word by word, which, while effective for capturing short-term dependencies, poses challenges in learning long-range dependencies.
RNNs encounter difficulties with long-range dependencies due to issues like vanishing or exploding gradients. This poses a limitation in capturing complex dependencies over long sequences. Architectures such as Long Short-Term Memory network (LSTM) (Hochreiter and Schmidhuber, 1997) and Gated Recurrent Units (GRUs) (Chung et al., 2014) were introduced to address these issues.
Choi et al. (Choi et al., 2016) used RNNs to generate stories by predicting the next sentence vector based on the current input, ensuring the continuity of the narrative flow within a story. Jain et al.(Jain et al., 2017) used GRUs to encode sequences of variable length input descriptions and to decode them as story-like summaries. Fig. 5(a) shows an example input provided to the model, and Fig. 5(b) shows an example story generated by the model. RNNs base their narratives generation decisions on statistical associations derived from the datasets on which they are trained, meaning that narrative elements are learned implicitly from the training data, leading to outputs that can sometimes be unpredictable. In the example we can see this unpredictability in the repetition of the phrase “we also saw some lily pads .”.
[…]
(a) Input captions.
[…]
(b) Generated story.
Figure 5. A story generated by an RNN (Jain et al., 2017).
4.6. Transformers and Large Language Models
Transformers (Vaswani et al., 2017) are a class of deep-learning models that represent a paradigm shift in sequential data processing and are the basis of the Large Language Models (LLMs) revolution. In contrast with RNNs, which process words sequentially in a recurrent way, transformers process the entire sentence simultaneously, eliminating the risk of “forgetting” past information and improving the training process efficiency.
LLMs are composed of multiple transformer layers, which enable them to attend to information at the word, sentence, and paragraph levels. This multi-layered approach allows for the construction of narratives that are not only coherent but also perceived as richly detailed and creatively varied. LLMs also leverage their extensive training data to produce stories across a wide range of styles and genres. However, as with RNNs, LLMs do not explicitly learn narrative elements or the underlying structure of storytelling. Their ability to generate stories originates from the statistical patterns and associations found within their training corpora. Nevertheless, the transformer architecture’s ability to attend to information across different parts and levels of the text enables LLMs to maintain context and coherence over diverse textual structures, contributing to their effectiveness in generating narrative content.
Bidirectional Encoder Representations from Transformers (BERT) (Devlin et al., 2019) is one of the first LLMs. It was widely adopted and established several state-of-the-art results for downstream tasks, even though it is now considered a small model: the base version has 110 million parameters, while current state-of-the-art models, such as Generative Pre-trained Transformer (GPT)-4 (OpenAI, 2024b) and Mixtral (team, 2023), have hundreds of billions or even trillions of parameters. Current LLMs, such as GPT (OpenAI, 2024b), have demonstrated proficiency in understanding context, generating coherent text, and even simulating human-like conversational abilities.
Some of the key aspects that characterize a LLM and are crucial for their effectiveness in narrative generation are their size (number of parameters), the nature of the training data, and the context window size. Regarding size, larger models tend to perform better, capturing more complex language nuances. However, increasing model size also requires more computational resources for training and inference. Regarding training data, LLMs benefit from diverse and extensive training datasets: a broader range of topics and writing styles in the training data helps the model generalize well to various tasks. The context window size refers to the number of preceding tokens the model considers when generating the next token. This is a critical feature, as it determines how much of the previous text the model can “remember” and use for making predictions about what comes next. A larger context window allows the model to understand longer dependencies in the text, improving coherence and context awareness. This is especially important for tasks that involve lengthy discussions or complex narratives. However, larger context windows come with increased computational costs and this relationship is quadratic (Vaswani et al., 2017), posing significant challenges in terms of memory and processing power, especially for very large models and has been a key focus of recent research efforts (Ding et al., 2023).
Several works using LLMs have been proposed for story generation. Tang et al. (Tang et al., 2022) proposed TextBox 2.0, a library that uses LLMs to generate stories. We do not include an example of a story generated by TextBox 2.0 as the authors did not provide one in their paper. Yang et al. (Yang et al., 2022) proposed Re3, a framework that prompts a general-purpose LLM to build a story plan consisting of a setting, characters, and a plot. This structured plan is generated from a short sentence describing the story, called premise. Story passages are then generated by repeatedly injecting contextual information from both the plan and current story state via prompt. This approach is one of the first approaches to explicitly model the elements character, setting, and plot in the story generation process. Fig. 6 shows the initial setup and outline for an example premise. The resulting story incorporates all the elements from the premise, including character interactions, setting descriptions, and plot developments. We did not include the generated story as it occupies several pages and is available in the original paper (Yang et al., 2022).
[…]
Premise: A woman discovers that her husband has been cheating on her and decides to get revenge.
Setting: The story is set in the present day, in a small town.
Characters:
1. Diane Chambers is a beautiful woman in her early thirties. She has long dark hair and blue eyes. She is married to Mark Chambers and is a stay-at-home mom.
2. Mark Chambers is a successful businessman in his early forties. He is tall and handsome, with brown hair and green eyes. He is married to Diane Chambers and is the father of two young children.
3. Karen Johnson is a beautiful blonde women in her early thirties. She is Mark Chambers’ mistress and is also married with two young children.
Outline:
1. Diane Chambers discovers that her husband has been cheating on her with another woman.
2. Diane Chambers decides to take revenge on her husband by having an affair of her own.
[…]
Figure 6. Re3 initial setup and outline for the shown premise (Yang et al., 2022).
4.7. Discussion
A critical limitation of rule-based, grammar-based, plan-based, and case-based methods is that they struggle with language variation, especially (but not only) with pronoun variation. This results in the frequent and repetitive use of certain structures and constituents, such as character names, leading to less fluid and less engaging narratives. The principle of “Show, Don’t Tell” (Dietrich and Sundell, 1983) is often underused: showing is more powerful than telling because it allows the reader to experience the story directly, rather than being told what to feel. However, these systems tend to rely on direct exposition to convey character emotions and story developments, rather than implying them through actions and interactions. This failure is not because the methods cannot solve these issues. Rather, they would require an impractically large number of conditions and rules to be manually encoded. RNNs are not immune to these challenges, but their inherent adaptability and capacity for learning from extensive datasets make them more capable of capturing the complexities of narrative storytelling. Likewise, LLMs are able to create fluid narratives with varied sentence structures, dialogues between characters, rich and descriptive details about the setting and characters, because they are able to learn aspects of the world from their training data. This allows them to generate details about the setting and characters that are not explicitly provided in the prompt. LLMs, however, are not without their limitations: they may suffer from hallucinations, that is, generating details that completely deviate from the original prompt, or generate contradictory or nonsensical information.
5. From NLP to computer vision
This section provides an overview of areas at the intersection of computer vision and Natural Language Processing (NLP) that are relevant to the topic of VSG, namely image and video captioning and Visual Question Answering (VQA). These tasks share a common challenge: they require the understanding of visual content and the ability to generate text based on that visual content. Some of the techniques used in these tasks have been adapted to VSG. Special attention is given to approaches based on deep learning architectures.
5.1. Image and Video Captioning
The transition from image and video captioning to VSG represents a shift from describing individual images or video segments to narrating image sequences or video segments. While the former focus on what is in images or videos, the latter also considers why and how, capturing the underlying narrative that links a series of images or video frames.
VSG models need to understand the temporal progression of events, causal relationships, and the ability to construct cohesive and engaging stories. Therefore, they build upon the foundational principles of image and video captioning, incorporating both the interpretation of visual data and the generation of coherent, contextually appropriate narratives.
The tasks of image and video captioning consist in generating descriptive, human-readable sentences that accurately represent the content of an image or video, identifying the main components in a visual scene and their relationships, alongside the ability to express these details in natural language. Initial efforts in image and video captioning relied on template-based and rule-based approaches. Current deep learning models are capable of end-to-end training, resulting in a substantial enhancement in the quality of the generated captions.
One of the most notable contributions in this domain was the application of encoder-decoder architectures, known as the ‘show and tell’ approach (Vinyals et al., 2014; Yao et al., 2015; Anderson et al., 2017). This technique employs Convolutional Neural Networks (CNNs) as encoders to extract visual features. Specifically, 2D-CNNs are employed for image processing where the model interprets two dimensions: height and width. On the other hand, for video processing, 3D-CNNs are typically used where an additional dimension, time or depth, is taken into account to perceive the temporal dynamics in video sequences. Then, RNNs or their improved variants, LSTMs (Hochreiter and Schmidhuber, 1997) and GRUs (Chung et al., 2014), known for their ability to effectively model sequential data, are commonly used as decoders in image and video captioning tasks, enabling the generation of coherent and contextually meaningful textual captions and descriptions from the visual features extracted by the CNNs. This combination of CNNs and RNNs effectively bridges the gap between vision and language.
In the realm of image captioning, transformers have been used as encoders to process visual features extracted from images. Leveraging the self-attention mechanism inherent in transformers, these models are able to capture relationships between visual elements, attending to relevant regions in the image, and highlighting visual cues and their interdependencies (Olimov et al., 2021; Radford et al., 2021; Liu et al., 2021a; Yu et al., 2019). Transformer models also demonstrate the ability to generate precise and elaborate textual descriptions, significantly improving the overall quality of image captions.
Similarly, transformers have made contributions to video captioning, addressing the challenges associated with understanding and describing temporal dynamics in videos. By treating video frames as a sequential series of images, transformers can leverage their attention mechanism to capture long-range dependencies across frames (Liu et al., 2021b; Lin et al., 2021) This enables the models to map the temporal progression of events, infer causal relationships, and grasp the underlying narrative structure within the video.
The integration of transformers in image and video captioning has yielded improvements that surpass the performance of traditional encoder-decoder architectures (Liu et al., 2021a, b). Transformers not only generate semantically accurate captions but also exhibit enhanced coherence and contextual relevance. Their attention mechanism allows for a incorporation of local and global context, enabling a more holistic understanding of the visual content and facilitating caption generation.
The use of transformers in image and video captioning holds great promise, but there are challenges to be addressed. Efficient fusion of visual features with textual information (Radford et al., 2021), handling long-term dependencies in video sequences (Liu et al., 2021b), and ensuring temporal coherence in caption generation for longer videos are among the ongoing research topics.
5.2. Visual Question Answering
Visual Question Answering (VQA) combines computer vision and natural language processing to enable the extraction of information from visual content, identifying objects, scenes, actions, and comprehending context, and provide meaningful answers to questions about that content. Both VQA and VSG involve understanding visual content and generating meaningful responses and the developments in VQA can be directly applied in VSG tasks. Through the use of VQA techniques, VSG models can gain valuable insights into the content and context of images and videos, improving the overall quality of their visual narratives.
Encoder-classifier deep-learning VQA models encode visual and textual inputs and select an answer from a predefined set using a classification process. Image encoders, typically pretrained CNNs, extract visual features from the input images. Question encoders, often RNNs, processes the input questions, converting them into fixed-length latent representations. Fusion techniques combine the visual and textual features, creating a fused representation that captures their interactions that is then classified using fully connected layers, mapping them to possible answer classes. Traditional VQA models are trained using supervised learning with large-scale datasets, where image-question-answer triplets are provided as training examples. The parameters are optimized to minimize the loss between predicted answers and ground truth answers during training. Some examples are BUTD (Anderson et al., 2017), that uses an attention-based model that combines bottom-up and top-down mechanisms to improve performance in image captioning and visual question answering tasks, and “Show, Ask, Attend, and Answer” (Kazemi and Elqursh, 2017), that combines LSTMs with multiple attention layers.
The transition from encoder-classifier methods to transformer methods in VQA was driven by the transformer’s ability to capture long-range dependencies and global contextual information. Transformers offered the potential to more efficiently capture the interactions between visual and textual modalities, enabling a more comprehensive understanding of the relationship between images and questions. In these VQA models, images and questions are encoded using transformer encoders and self-attention mechanisms capture the dependencies between words and the visual features, allowing the model to focus on relevant image regions while processing the question. The fused encoded representations are then fed into classification layers to generate answers (Bao et al., 2022; Wang et al., 2023a, 2022, b). Recent advances in VQA also explore new models that approach the task as a generation problem rather than a classification one, allowing for a more flexible and diverse range of answers (Chen et al., 2022).
6. Visual Story Generation
The current state-of-the-art and baseline methods employed in VSG have evolved from their origins in textual story generation. These methods have served as foundations for the development of more sophisticated techniques and continue to provide a comparative baseline for new models and approaches.
6.1. Visual Storytelling with Convolutional and Recurrent Neural Networks
Huang et al. (Huang et al., 2016) proposed a dataset of sequential vision-to-language mappings, SIND v.1 (later renamed to VIST), with 81,743 images in 20,211 sequences, with both descriptive (captions) and story language, and aiming at facilitating research, from understanding individual images to comprehending image sequences that narrate events over time.
This work also proposed models for generating stories based on image sequences, using a sequence-to-sequence RNN approach, which extends the image captioning technique. It extracts features from each image using a pretrain CNN called AlexNET (Krizhevsky et al., 2012). It then encodes image sequences by using a GRU over the features of each individual image. The story decoder model consists of another GRU and generates narratives word by word. Four different decoding methods, shown in Table 3, were considered: beam search, greedy search, greedy search with removal of duplicates, and greedy search with removal of duplicates and incorporation of grounded words.
Decoding Method METEOR
Beam=10 23.13
Greedy (Beam=1) 27.76
Greedy - Duplicates 30.11
Greedy - Duplicates + Grounded 31.42
Table 3. Results of each decoding approach using evaluated using METEOR scores on the VIST (Huang et al., 2016) dataset.
Results indicate that beam search alone does not perform well in generating high-quality stories. This contrasts with the results for image captioning, where beam search is more effective than greedy search. The reason is that generating stories with beam search results in generic high-level descriptions such as “This is a picture of a dog” originating from the label bias problem. On the other hand, by using a beam size of 1 (greedy search) there is an improvement in the quality of the generated stories.
Since stories generated using a greedy search may sometimes contain repeated words or phrases that can adversely affect textual quality, a duplicate-removal heuristic was added to ensure that content words cannot be produced more than once on a given story. Finally, to further improve quality, the previous approach was extended to include “visually grounded” words. These are words that are authorized for use by the story model only if they are licensed by the caption model: for a word to be included in the story, it must have appeared with some frequency in the captions generated by the caption model for the sequence of images. Fig. 7 shows an example of the generated stories for an image sequence.
Figure 7. Sequence of images used to generate stories using different decoding methods. Beam search: This is a picture of a family. This is a picture of a cake. This is a picture of a dog. This is a picture of a beach. This is a picture of a beach. Greedy: The family gathered together for a meal. The food was delicious. The dog was excited to be there. The dog was enjoying the water. The dog was happy to be in the water. Greedy - duplicates: The family gathered together for a meal. The food was delicious. The dog was excited to be there. The kids were playing in the water. The boat was a little too much to drink. Greedy - duplicates + grounded words: The family got together for a cookout. They had a lot of delicious food. The dog was happy to be there. They had a great time on the beach. They even had a swim in the water. Adapted from (Huang et al., 2016).
6.2. Visual Storytelling with Convolutional and Bidirectional Recurrent Networks Networks
Park et al. (Park and Kim, 2015) proposed Coherent Recurrent Convolutional Network (CRCN), integrating various neural network components to bridge the gap between visual content and textual narrative: VGGNet (Simonyan and Zisserman, 2015) as image encoder and Bidirectional Recurrent Networks (BRNN) to capture contextual information from both the forward and backward directions within a given text sequence. CRCN also introduces a local coherence model that focuses on maintaining coherence and context within the generated sentence sequences. It also utilizes parse trees to detect grammatical roles and structural attributes, ensuring that the generated sentences integrate with the broader narrative context. During training, correctly aligned image-sentence pairs are given higher scores than misaligned pairs.
The evaluation of CRCN considered several automatic metrics and several baseline works, including one that excludes the entity coherence model. Table 4 shows the results for the NY City (Park and Kim, 2015) and Disneyland (Park and Kim, 2015) datasets. Both datasets were introduced by Park et al. (Park and Kim, 2015) and consist of image sequences with the corresponding blog posts. In these datasets, results were superior to all baselines. The exception, where it was a close second, was with the baseline without entity coherence and in only certain metrics. The work was further validated by user studies conducted via Amazon Mechanical Turk (AMT), showing that human evaluators prefer CRCN-generated sequences over several baselines.
Metric B-1 B-2 B-3 B-4 C M R@1 R@5 R@10
NY City 26.83 5.37 2.57 2.08 30.9 7.69 11.67 31.19 43.57
Disneyland 28.40 6.88 4.11 3.49 52.7 8.78 14.29 31.29 43.20
Table 4. Performance Metrics for CRCN on New York City and Disneyland Datasets. Bold values represent state-of-the-art results at the time. B stands for BLEU, C for CIDEr, M for METEOR and R for Recall (Park and Kim, 2015).
Fig. 8 shows an example of a sequence of images and the corresponding generated sentences.
Figure 8. Sequence of images used by CRCN a story. (1) One of the hallway arches inside of the library (2) As we walked through the library I noticed an exhibit called lunch hour nyc it captured my attention as I had also taken a tour of nyc food carts during my trip (3) Here is the top of the Chrysler building everyone’s favorite skyscraper in new york. (4) After leaving the nypl we walked along 42nd st. (5) We walked down fifth avenue from rockefeller centre checking out the windows in saks the designer stores and eventually making our way to the impressive new york public library. Adapted from (Park and Kim, 2015).
Besides, Park et al. (Park and Kim, 2015) which is the first approach to use a bidirectional RNN for visual storytelling, to the best of our knowledge, XE (Wang et al., 2018) and AREL (Wang et al., 2018) have also proposed models for visual storytelling that use bidirectional RNNs on the VIST dataset.
6.3. Transitional Adaptation of Pretrained Models for Visual Storytelling
Yu et al. (Yu et al., 2021) proposed Transitional Adaptation of Pretrained Model (TAPM), aimed at refining the generation of textual descriptions for visual content, particularly in the context of VSG tasks. It aims at bridging the gap between pretrained language models and visual encoders. In contrast with previous models for vision-to-language generation tasks, which typically pretrain a visual encoder and a language generator separately and then jointly fine-tune them for the target task, TAPM proposes a transitional adaptation task to harmonize the visual encoder and language model for downstream tasks like VSG, mitigating discord between visual specificity and language fluency arising from separate training on large corpora of visual and text data. TAPM introduces an approach that adapts multimodal modules through a simpler alignment task focusing solely on visual inputs, eliminating the need for text labels.
TAPM components aim at enhancing the quality of textual descriptions for visual content in storytelling tasks. They are a visual encoder, a language generator, adaptation training, sequential coherence loss, training with adaptation loss, and a fine-tuning and inference process. The visual encoder, a pretrained model, extracts features from images or videos. In TAPM, it becomes integral during the adaptation phase, where it integrates with the language generator to fuse visual and textual information. The pretrained language generator model is responsible for converting visual information into textual descriptions. During the adaptation phase, it generates video and text embeddings, aligning textual representations with the corresponding visual features, based on a sequential coherence loss function.
The loss function divides sequential coherence into three components: past, current, and future matching losses. The past matching loss uses a Fully Connected (FC) layer $f_{p}$ to project the text representation $b_{s_{i}}$ of video $i$, drawing it nearer to the visual representation $v_{b_{i-1}}$ of the preceding video $i-1$ and distancing it from those of non-sequential videos. The future matching loss projects $b_{s_{i}}$ via a distinct FC layer $f_{f}$, aligns with the subsequent visual representation $v_{b_{i+1}}$. The current matching loss then aligns the current visual representation $v_{b_{i}}$ with $b_{s_{i}}$ through another FC layer $f_{c}$. These components are unified by the FC layer projections in their respective visual spaces, pulling the embeddings of correct matches closer and pushing incorrect matches further apart. Margin ranking losses are utilized to implement this concept, contrasting correct matches with incorrect ones. The final sequential coherence loss for a given video $i$ is formulated as shown in Eq. 15, where $\cos$ is the cosine similarity, and $j$ represents indices of incorrect matches.
(15) $L_{i}=$ $\sum_{j\neq i-1}\max(0,1+\cos(v_{b_{j}},f_{p}(b_{s_{i}}))-\cos(v_{b_{i-1}},f_{p}(b_{s_{i}})))$
$+$ $\sum_{j\neq i}\max(0,1+\cos(v_{b_{j}},f_{c}(b_{s_{i}}))-\cos(v_{b_{i}},f_{c}(b_{s_{i}})))$
$+$ $\sum_{j\neq i+1}\max(0,1+\cos(v_{b_{j}},f_{f}(b_{s_{i}}))-\cos(v_{b_{i+1}},f_{f}(b_{s_{i}}))),$
TAPM uses a split-training strategy to optimize model performance. Initially, the visual encoder undergoes adaptation training with the adaptation loss, while the text encoder and language generator remain fixed. Subsequently, all components are jointly updated with the generation loss, allowing the model to optimize the adaptation task before addressing the more challenging generation objective. After the adaptation and split-training phases, TAPM undergoes fine-tuning. The model is then ready for the inference phase, generating captions for unseen visual inputs.
Fig. 9(a) shows results for TAPM against selected baselines on the LSMDC 2019 and VIST datasets. Fig. 9(b) and 9(c) show human evaluation results in which TAPM surpasses adversarial baselines on LSMDC 2019. On VIST it surpasses the XE (Wang et al., 2018) and AREL (Wang et al., 2018) baselines in relevance, expressiveness, and concreteness. These results highlight the strengths in word choice and contextual accuracy, showcasing its ability to capture causal relationships between images. However, the score is still far from human performance, indicating that there is still room for improvement.
Models Public Test Blind Test
C R C R
Official Baseline (Park et al., 2018) 7.0 12.0 6.9 11.9
XE (Wang et al., 2018) 7.2 11.5 - -
AREL (Wang et al., 2018) 7.3 11.4 - -
TAPM 10.0 12.3 8.8 12.4
(a) Results on LSMDC 2019 public and blind test set. C stands for CIDEr and R for ROUGE.
Models Scores
Human 1.085
Official Baseline (Park et al., 2018) 4.015
TAPM 3.670
(b) Human evaluation results on the LSMDC 2019 blind test set according to a Likert scale from 5 (worst) to 1 (best), where lower is better.
TAPM vs XE TAPM vs AREL
Choice (%) TAPM XE Tie TAPM AREL Tie
Relevance 59.9 34.1 6.0 61.3 32.8 5.9
Expressiveness 57.3 32.3 10.4 57.3 34.0 8.7
Concreteness 59.1 30.3 10.7 59.6 30.4 10.0
(c) Human evaluation results on VIST. Higher is better.
Figure 9. TAPM evaluation results. Adapted from (Yu et al., 2021).
6.4. Combining LLMs (GPT-2) and Character Features
Hong et al. (Hong et al., 2023) proposed CharGrid (Character-Grid Transformer), a transformer model that integrates diverse input features, from global image attributes to character-specific details. It takes a sequence of input tokens, including global image features obtained using Swin Transformer’s (Liu et al., 2021a), character features extracted by cropping character instances, a character grid for coherence assessment, and text features introduced incrementally during processing. The trainable image and character encoders process the image feature inputs. The character grid is flattened and fed to a FC layer. The transformer module then processes these inputs, generating an output as a probability distribution over potential tokens using a pre-trained GPT-2 tokenizer. The model architecture is shown in Fig. 10.
Figure 10. CharGrid architecture: the blue boxes are pre-trained components where the parameters are fixed. Adapted from (Hong et al., 2023).
During training, model parameters, excluding the vision backbone, are randomly initialized. The training objective involves maximizing the likelihood of image sequence-story pairs through backpropagation. Nucleus sampling (Holtzman et al., 2019) is employed for validation, and the METEOR score is used for evaluation.
CharGrid has two additional variants: ObjGrid (Object-Grid Transformer) and EntiGrid (Entity-Grid Transformer). ObjGrid replaces character features with object features. This model incorporates global features and object features, with an object grid representing coherence based on the similarity between global image features and object features. It is similar to CharGrid but includes both character and object features. EntiGrid incorporates global, character, and object features, with an entity grid representing coherence through the similarity between global image features and both character and object features. These variants explore the impact of different input feature combinations on model performance. CharGrid, with its emphasis on character coherence, serves as the primary model, while ObjGrid and EntiGrid address the contributions of object-related features.
The results in Table 5 show that CharGrid outperforms TAPM with character features and GPT-2 with character features on various metrics, emphasizing the effectiveness of character grid representations for coherence in VSG. A crowd-sourcing experiment was conducted with 28 workers to obtain binary judgments on grammaticality, coherence, diversity, and visual groundedness of the generated stories. The results in Table 6 show that TAPM with character features excels in visual groundedness over plain TAPM, while CharGrid surpasses TAPM with character features across all metrics. Two-sided binomial tests support that the character grid representation yields superior stories, affirming the reference-based metrics findings.
Model B-1 B-2 B-3 B-4 M R-L C
GPT-2 38.65 20.28 9.78 4.68 31.64 24.24 1.66
GPT-2 + obj 40.65 21.35 10.2 4.87 31.69 24.05 1.85
GPT-2 + char 39.95 21.04 10.11 4.92 31.85 24.19 1.57
GPT-2 + obj,char 40.41 21.44 10.56 5.06 32.03 24.38 1.87
TAPM 39.85 21.7 10.72 5.19 32.38 25.09 1.48
TAPM + obj 40.86 22.13 10.83 5.25 32.34 24.91 1.82
TAPM + char 40.03 21.68 10.66 5.18 32.42 24.88 1.4
TAPM + obj,char 40.87 21.99 10.72 5.06 32.48 24.87 1.59
ObjGrid + obj 47.66 25.26 11.95 5.42 32.83 24.42 4.68
EntityGrid + obj,char 45.83 24.85 12.11 5.7 32.68 24.89 3.53
CharGrid + char 47.71 25.33 11.95 5.42 33.03 25.01 4.83
Table 5. Results using different input features on the test set of VWP. Global features are always included. “+ obj” and “+ char” represent that object and character features are also included, respectively. BLEU (B), METEOR (M), ROUGE-L (R-L), and CIDEr (C) values are averages of three runs with different random seeds. Adapted from (Hong et al., 2023).
Model Grammaticality Coherence Visual Groundedness Diversity
TAPM + char vs. TAPM +2.45 +1.99 +3.99* +1.69
CharGrid vs. TAPM + char +6.49** +8.41** +6.25* +11.06**
Table 6. Human binary judgments in percentage of generated stories between TAPM and TAPM with character features (TAPM + char), TAPM + char and CharGrid on the test set of VWP across four criteria: Grammaticality, Coherence, Visually Groundedness, and Diversity. The numbers are percentages. * means p-value $<$ 0.05. ** means p-value $<$ 0.01. Adapted from (Hong et al., 2023).
6.5. Visual Story Generation using Graphs for Event Ratiocination
Zheng et al. (Zheng et al., 2021) introduced Hypergraph-Enhanced Graph Reasoning (HEGR) for “visual event ratiocination”, a task that involves generating interpretations and narratives for events that precede, coincide with, or follow a visual scene, as well as discerning the intents of depicted characters. The work addresses several downstream tasks, including VSG. The work introduces a framework to improve the integration and interpretation of visual and textual data, using hypergraphs to address challenges in multimodal interactions and temporal dynamics. Figure 11 shows this architecture.
Figure 11. Architecture of HEGR by Zheng et al. Adapted from (Zheng et al., 2021).
The method begins with feature extraction. Visual features are extracted using Mask R-CNN (He et al., 2017) to detect objects and their spatial characteristics within images. Textual features are extracted from associated descriptions or captions using BERT (Devlin et al., 2019), capturing semantic information. In the second phase, semantic graphs are constructed for each modality: nodes represent detected features or semantic tokens, while edges depict relationships within the same modality. To capture dependencies between modalities and across various temporal segments, hypergraphs are employed, linking semantic graphs based on contextual and temporal relevance, forming a high-dimensional interaction map. In the third phase, graph convolution is used to update node representations within each graph, allowing contextual information aggregation through learned adjacency matrices. Hyper-graph convolution extends this to include higher-order relationships across multiple graphs and temporal instances. In the fourth phase, a self-attention mechanism refines graph representations by dynamically adjusting the significance of nodes within and across modalities. This process focuses on the most relevant features for narrative generation. Finally, the event ratiocination generator uses the enriched feature representations to generate narratives using BART (Lewis et al., 2019).
The method was evaluated on the VIST dataset, achieving state-of-the-art performance in terms of BLEU-4, ROUGE-L and METEOR scores achieving 16.4, 37.4, and 37.8, respectively.
6.6. Discussion
The works presented in this section have made substantial contributions to the field of VSG, establishing robust benchmarks for the assessment of emerging models and methodologies.
Specifically, the first three techniques, presented in Sections 6.16.2, and 6.3 employed a CNN encoder for visual feature extraction and a RNN decoder for language generation. The methods detailed in Sections 6.1 and 6.3, offer comparative analyses on the VIST dataset, with METEOR scores of 31.4 and 37.2, respectively. Notably, the latter technique surpasses the former, which can be attributed to its transitional training approach.
The study outlined in Section 6.4, CharGrid, does not present findings on the VIST dataset. However, this method is benchmarked against the method detailed in Section 6.3, TAPM, on the VWP dataset, where it demonstrates superior performance across all evaluated metrics. This comparison is shown in Table 5, specifically within the rows labeled “TAPM + obj,char” and “CharGrid”. This method differentiates itself from the previous by focusing on the extraction of object-level features from visual inputs rather than employing a CNN to encode the entire visual input. Furthermore, it leverages the capabilities of an LLM, using its extensive world knowledge to enhance language modeling, illustrating a novel integration of visual and linguistic elements in story generation.
Finally, the method discussed in Section 6.5, HEGR, introduces a different approach that uses hypergraphs to model the contextual and temporal relationships between visual and textual data. This approach is explicitly able to model some story elements, namely the characters and the setting.
Table 7 summarizes the performance of the methods described above, along with other methods from the literature that contributed to the field of VSG. The listed models range from earlier methods using CNNs and RNNs to more recent advances that integrate object detection models like Faster R-CNN (Ren et al., 2015) and Mask-RCNN (He et al., 2017), and LLMs like GPT-2 (Radford et al., 2019) and BART (Lewis et al., 2019). The first two methods, Park et al. (2015) (Park and Kim, 2015) and Park et al. (2019) (Park et al., 2018) are not comparable to the other methods as they do not use the VIST dataset. The remaining methods are evaluated on the VIST dataset with the exception of CharGrid (2023) (Hong et al., 2023) that is evaluated on the VWP dataset. The authors of CharGrid trained and evaluated the performance of TAPM on the VWP (Hong et al., 2023) dataset, making the two methods directly comparable and it is likely that CharGrid would perform slightly better than TAPM on the VIST dataset. It is however not possible to know if CharGrid would perform better than HEGR on the VIST dataset without further experiments.
Model Encoding Decoding Dataset B-4 R-L M
Park et al. (2015) (Park and Kim, 2015) CNN, Bi-RNN RNN Blog Posts 3.49 - 8.78
Park et al. (2019) (Park et al., 2018) Resnet-101/152, LSTM LSTM ActivityNet 9.91 - 16.48
Huang et al. (2015) (Huang et al., 2016) CNN, GRU GRU VIST - - 31.4
Xu et al. (2015) (Xu et al., 2015) VGGNet LSTM VIST - 28.94 32.98
Yu et al. (2017) (Yu et al., 2017) ResNet-101, Bi-GRU GRU VIST - 29.53 34.12
CST (2018) (Gonzalez-Rico and Pineda, 2018) CNN, LSTM LSTM VIST 12.7 29.2 34.4
AREL (2018) (Wang et al., 2018) ResNet-152, Bi-GRU GRU VIST 14.1 29.5 35.0
HSRL (2019) (Huang et al., 2018) ResNet-152 LSTM VIST 12.32 30.84 35.23
GLACNet (2019) (Kim et al., 2018) ResNet-152, Bi-LSTM LSTM VIST - - 30.63
StoryAnchor (2020) (Modi and Parde, 2019) ResNet-152, Bi-GRU GRU VIST 14.0 30.0 35.5
INET (2020) (Jung et al., 2020) CNN, GRU GRU VIST 14.7 29.7 35.5
AOG-LSTM (2023) (Liu et al., 2023) RestNet-152 LSTM VIST 12.9 30.1 36.0
CoVS (2023) (Gu et al., 2023) RestNet-152, GRU GRU VIST 15.2 30.8 36.5
TAPM (2021) (Yu et al., 2021) 1 Faster R-CNN, ResNet-152 GPT-2 VIST - 33.1 37.2
HEGR (2021) (Zheng et al., 2021) 1,2 Mask-RCNN, BERT BART VIST 16.4 37.4 37.8
TAPM (2021) (Hong et al., 2023) Faster R-CNN, ResNet-152 GPT-2 VWP 5.19 25.09 32.38
CharGrid (2023) (Hong et al., 2023)2 Mask-RCNN, SwinTransformer GPT-2 VWP 5.42 25.01 33.03
Table 7. Analysis of Visual Story Generation works. The columns “B-4”, “R-L”, and “M” represent the BLEU-4, ROUGE-L, and METEOR scores, respectively. All works, except Huang et al. (2015) (Huang et al., 2016) and CST (2018) (Gonzalez-Rico and Pineda, 2018), use attention mechanisms. All works use images as input, except ActivityNet (Fabian Caba Heilbron and Niebles, 2015) (it uses videos). The mark 1 indicates works that used extra training data (direct comparison may not be fair). The mark 2 indicated works that explicitly model some story elements, namely the characters by the detection of objects in the case of CharGrid (2023) and HEGR (2021) and the setting in the case of HEGR (2021).
7. Conclusions
This survey reviewed the field of story generation, explaining its development and current methods. It started by explaining basic story elements and early techniques, which mostly involved manually created rules. We looked at the datasets and evaluation metrics used to judge the quality of the stories. The discussion pointed out the need for better metrics to assess complex tasks like story generation accurately. The document also covered related areas like image and video captioning and VQA, discussing how improvements in these fields have helped enhance VSG.
The survey also looked at the latest developments in VSG and discussed various models that use deep learning. It started with basic models using simple encoder-decoder setups and went on to more advanced systems that use transformers. Each model was described with a focus on how they handle issues like making the story coherent, keeping it relevant to the context, and combining detailed visual and textual information.
Despite the advances, challenges remain for VSG. In the following subsections, we discuss some of the key challenges and opportunities in the field that could guide future research directions. We start with the ones that are more concrete and could influence the field in the short term and end with the ones that are more abstract and could potentially influence the field in the long term.
7.1. Exploring Large Language Models
The field of VSG using LLMs is under-explored, with only a few works delving into this domain. To the best of our knowledge the existing works predominantly rely on GPT-2, which is an outdated LLM when compared to models such as Mixtral (team, 2023) which is the best performing open-source model on MT-Bench or GPT-4 which is the best performing proprietary model. The scarcity of the integration of LLMs with visual elements presents an opportunity for further research, exploring the potential of these models in generating more coherent and contextually relevant stories.
7.2. Automatic Evaluation Metrics
The need for robust and efficient automatic evaluation metrics is emphasized in our survey. While existing metrics like METEOR and CIDEr have been widely adopted, they do not fully capture the nuances of story quality. The metrics mostly focus on the similarity between the generated text and the reference text. A story can be retold in various ways, employing different perspectives while preserving the original message. Consequently, a generated text might deviate significantly from the reference text in terms of specific wording and still deliver the same story in a coherent way. Thus, while these metrics provide valuable insights into certain dimensions of text generation, they did not fully encompass the multifaceted nature of storytelling. As LLMs continue to improve, they show potential in replacing human annotators in many tasks (Gilardi et al., 2023) and so metrics to evaluate stories could be developed based on LLMs. We advance that these metrics could be based on the existing systematic scoring systems (Zheng et al., 2023) that were developed to evaluate the performance of LLM-based chat bots.
7.3. Decomposition of Story Generation
Addressing the complexity of stories, we underscore the importance of decomposing the story generation process into modular components. Each module can focus on specific attributes such as author goals, character interactions, consistency, etc. This decomposition has already been explored with modeling characters in CharGrid (Yu et al., 2021) and the modeling of characters and setting in HEGR (Zheng et al., 2021). We believe that explicit modeling of other story elements and abstract concepts such as emotions and culture, possibly using hypergraphs or similar structures could further enhance the quality of generated stories.
7.4. Hybrid Systems and Knowledge Integration
Building on the decomposition concept, hybrid systems, combining planning-based techniques and machine learning approaches, present an intriguing direction. Hyper-graphs, as seen in HEGR (Zheng et al., 2021), could possibly be a starting point as a deep learning approach to model the planning process replacing the traditional handcrafted planning systems. The integration of multiple knowledge sources can enrich the diversity and quality of generated stories, paving the way for more robust and context-aware narrative creation.
Acknowledgements.
Daniel A. P. Oliveira is supported by a scholarship granted by Fundação para a Ciência e Tecnologia (FCT), with reference 2021.06750.BD. Additionally, this work was supported by Portuguese national funds through FCT, with reference UIDB/50021/2020.

### v2

@@
-Story Generation from Visual Inputs: Techniques, Related Tasks, and Challenges
-Abstract.
+Visual Story Generation; Image Captioning; Visual Question Answering; Storytelling
+Abstract
 Creating engaging narratives from visual data is crucial for automated digital media consumption, assistive technologies, and interactive entertainment.
 This survey covers methodologies used in the generation of these narratives, focusing on their principles, strengths, and limitations.
-The survey also covers tasks related to automatic story generation, such as image and video captioning, and visual question answering, as well as story generation without visual inputs.
-These tasks share common challenges with visual story generation and have served as inspiration for the techniques used in the field.
+The survey also covers tasks related to automatic story generation, such as image and video captioning, and visual question answering.
+These tasks share common challenges with Visual Story Generation (VSG) and have served as inspiration for the techniques used in the field.
 We analyze the main datasets and evaluation metrics, providing a critical perspective on their limitations.
-Visual Story Generation, Image Captioning, Visual Question Answering, Storytelling
-copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXjournal: CSURjournalvolume: 00journalnumber: 0article: 000publicationmonth: 0ccs: Computing methodologies Natural language generationccs: Computing methodologies Video summarizationccs: Computing methodologies Scene understanding
-1. Introduction
-Stories are fundamental to human experience, serving as a bridge between imagination and reality (Hühn et al., 2014).
+keywords:
+Visual Story Generation; Image Captioning; Visual Question Answering; Storytelling
+[…]
+[…]
+1 Introduction
+In this document we survey the field of story generation from visual inputs, covering techniques, related tasks, and challenges.
+Stories are fundamental to human experience, serving as a bridge between imagination and reality Hühn et al.
+(2014).
 They represent the art of telling tales, whether real or imagined, and present a sequence of events that can captivate, inform, and provoke thought.
@@
 Stories have played a crucial role in shaping societies, passing down knowledge and values from one generation to another.
-Whether told around a campfire, written in books, or shared through videos and interactive platforms, storytelling continues to captivate and connect people, transcending time and technology (Abbott, 2008).
-The field of automated story generation has gained attention, fueled by advances in Natural Language Processing (NLP) (Jurafsky and Martin, 2009) and computer vision (Szeliski, 2010).
-The goal is to create systems capable of generating novel and engaging stories, similar to those produced by humans (Gervas, 2009).
-With the increasing availability of images and videos, there is a growing interest in generating stories based on this visual data (Krishna et al., 2017).
-We refer to this task as Visual Story Generation (VSG).
-1.1. Definition and challenges
+Whether told around a campfire, written in books, or shared through videos and interactive platforms, storytelling continues to captivate and connect people, transcending time and technology Abbott (2008).
+The field of story generation has gained attention, fueled by advances in Natural Language Processing (NLP) Jurafsky and Martin (2009) and computer vision Szeliski (2010).
+The goal is to create systems capable of generating novel and engaging stories, similar to those produced by humans Gervas (2009).
+With the increasing availability of images and videos, there is a growing interest in generating stories based on this visual data Krishna et al.
+(2017).
+We refer to this task as VSG.
+1.1 Definition and challenges
 VSG is the task of automatically creating coherent and engaging stories grounded on visual inputs, such as images or videos.
@@
 VSG serves as a robust indicator of machine comprehension of visual content and poses several challenges, some of which are:
-  • Understanding visual content: The recognition of objects and scenes in the visual data, as well as understanding the relationships between them is crucial for generating coherent and meaningful narratives (Xu et al., 2015).
+  • Understanding visual content: The recognition of objects and scenes in the visual data, as well as understanding the relationships between them is crucial for generating coherent and meaningful narratives Xu et al.
+(2015).
   • Temporal coherence: A story is inherently sequential and maintains a temporal order.
@@
 However, these metrics are not appropriate to capture the creativity and engagement aspects of a story.
-1.2. Background and Motivation
-The motivations for VSG ] are numerous.
+1.2 Background and Motivation
+The motivations for VSG are numerous.
 For example, VSG can provide an interactive component to media consumption.
@@
 By developing models that emulate human-like storytelling abilities, researchers can gain insights into the cognitive processes involved in story creation, as well as explore the underlying mechanisms of creativity.
-These approaches can further lead to novel applications, such as more engaging virtual experiences (Fan et al., 2018; Peng et al., 2018).
+These approaches can further lead to novel applications, such as more engaging virtual experiences Fan et al.
+(2018); Peng et al.
+(2018).
 These use cases represent only a fraction of the potential applications of VSG.
 As research in this field progresses, we anticipate that more applications will emerge, spanning an even broader range of domains.
-1.3. Scope and structure of the survey
-We start by defining the core elements of a story: they are essential for understanding the structure and dynamics of a story and served as foundational components for the development of the early story generation systems.
+1.3 Scope and structure of the survey
+We start by defining the core elements of a story in Section 2 as they are essential for understanding the structure and dynamics of a story.
 Section 2 presents these elements and discusses their significance in storytelling.
 It also clarifies the distinctions between story, narrative, and plot, illustrating how a single story can be narrated in multiple ways to yield diverse experiences.
-We then explore the main datasets used in story generation, and the various metrics used to assess their performance.
+In Section 3 we then explore the main datasets used in story generation, and the various metrics used to assess their performance.
 We address the limitations that current evaluation metrics face in such an open-ended task and discuss potential directions for future research in this regard.
-Section 3 addresses these issues.
-Since they form the background for VSG approaches, we also explore the methods for textual story generation: from the early rule-based systems to deep learning models.
-We focus on deep learning methodologies, especially transformer architectures and Large Language Models (LLMs) (Bolaños et al., 2015; Ammanabrolu et al., 2019), due to their state-of-the-art performance.
-Section 4 presents these approaches and emphasizes their contributions and historical significance.
 The connection of VSG with computer vision tasks such as Visual Question Answering (VQA) and image/video captioning is also acknowledged, since they share core challenges, such as visual understanding and narrative generation.
@@
 Section 5 presents a broad perspective on the state-of-the-art of VSG.
-Finally, Section 7 presents the key findings of this survey and proposes directions for future research.
-2. The Elements of Stories
-The essence of a story lies in its ability to connect a series of events and experiences, often conveyed by a narrator to an audience.
-These can range from personal intimate stories to larger epics that span cultures and time.
-A story can be told in multiple ways using different narratives.
-These narratives are more than mere descriptions or dramatic enactments.
-They involve a process of narration where events are selected, arranged, and em-plotted to give meaning to experiences.
-This structuring of events is what forms the plot or the story’s synopsis.
-The terms story, narrative, and plot are often used interchangeably in everyday language, but they have distinct meanings.
-According to the Oxford English Dictionary (Simpson, 2002) a story is the straightforward, chronological sequence of events, encompassing characters, settings, and actions.
-The story is the “what” – what happened, who it happened to, where and when it happened.
-In contrast, the narrative shapes the form or the “how”, adding context, interpretation, perspective, tone, and style of the storytelling, and nuance to the same set of events.
-The plot is connected to the narrative, referring to the sequence of events as they are presented to the audience.
-Thus, a single story can be narrated in multiple ways, each having a different plot and offering a distinct experience to the audience.
-Narratives can be broadly categorized as linear or nonlinear (Lau and Chen, 2008), depending on their structure.
-Each type offers a different approach to unfolding a story and has distinct implications for how the audience experiences and interprets it.
-Linear narratives follow a straightforward chronological progression, where events unfold in a sequential and cause-effect manner.
-Linear narratives often have a single protagonist or a group of characters whose actions and decisions drive the plot forward.
-This structure is commonly used in traditional storytelling formats.
-Linear narratives provide a sense of progression and closure, making them easy to follow and understand.
-Nonlinear narratives deviate from the traditional linear structure by presenting events out of chronological order.
-They often involve flashbacks, flash-forwards, or parallel storylines that disrupt the temporal sequence.
-These narratives can create intrigue, suspense, and complexity by challenging the audience’s expectations and requiring them to piece together the fragmented storyline.
-The film “Cloud Atlas” (Mitchell, 2008) is an example.
-Story elements – character, conflict, theme, setting, plot, and mode – are the essential components that form the foundation of a story and are vital for both crafting and analyzing stories.
-In the following subsections, we detail each of these elements, discussing their significance in shaping the story.
-2.1. Character
-Characters are the entities through which the story unfolds (DiBattista, 2011).
-Their decisions and actions are pivotal in propelling the plot forward.
-Characters are typically people or other beings, fictional or based on a real-life person, and through their actions and speech, they often convey essential motives and emotions.
-The way an audience initially perceives a character significantly influences their ongoing perception and emotional connection, such as empathy or antipathy.
-This initial impression sets the stage for the audience’s expectations regarding the future behavior of the character.
-Characters who deviate unexpectedly from their established patterns of behavior can create confusion or surprise, challenging the audience’s understanding and expectations.
-Narratives usually center around main characters or protagonists.
-They are the focal points of the story, often encountering central conflicts, undergoing significant personal growth, or acquiring new knowledge.
-Many stories also have antagonists, adding complexity and tension to the storyline.
-In traditional stories, the protagonist often takes the form of a hero, a sympathetic figure battling for morally good causes.
-Conversely, the antagonist may be portrayed as a villain, acting against these good causes or even perpetrating evil.
-Narratives often employ a diverse range of character types, each contributing uniquely to the story’s depth and dynamics.
-The way characters are developed, presented, and evolve throughout the narrative plays a crucial role in engaging the audience and driving the story’s progression.
-2.2. Conflict
-Conflict is a fundamental element that generates tension and drives the thoughts and actions of characters (Abbott, 2008).
-The manner in which conflicts are introduced, developed, and resolved plays a significant role in shaping the trajectory of the narrative and the engagement of the audience with the story.
-A well-crafted conflict not only propels the narrative forward but also enriches the emotional and psychological depth of the characters and the story itself.
-The resolution of the conflict, typically towards the end of the narrative, is a crucial aspect.
-It marks the culmination of tension and provides closure to the story’s central problem.
-Broadly, conflict encompasses any form of tension that influences characters, while more narrowly, it may refer to the main challenge faced by the protagonist of the story.
-This primary conflict often leads to additional struggles for the protagonist, such as feelings of anxiety, insecurity, indecisiveness, or other mental difficulties.
-These struggles can be categorized as secondary or internal conflicts, adding layers to the journey of the character and depth to the story.
-Longer stories typically involve multiple conflicts.
-These can be smaller-scale issues that occur alongside or as part of the main conflict, contributing to the complexity of the story.
-Commonly used conflict types include: “Character vs.
-Character” (e.g., opposition); “Character vs.
-Nature” (e.g., natural forces); “Character vs.
-Society” (e.g., societal norms); “Character vs.
-Unavoidable Circumstances” (e.g., fate); and “Character vs.
-Self” (e.g., dilemmas).
-2.3. Theme
-Themes encompass the major underlying ideas presented by a story, often abstract and open to the audience’s interpretation (Griffith, 2010).
-Unlike more concrete story elements, themes invite discussion and contemplation, allowing for a range of interpretations and conclusions.
-They are not definitive statements but rather exploratory concepts that encourage the audience to engage with and reflect on the story’s deeper meanings.
-As the narrative unfolds, themes can evolve or reveal new dimensions, offering the audience opportunity to develop their own ideas about the messages or insights the story conveys.
-The subjective nature of themes means that different audience members may arrive at varying interpretations, which might differ from or even go beyond the creator’s original intent.
-Thus, understanding and incorporating themes in narrative generation allows for a dynamic, thought-provoking storytelling experience that resonates with and challenges the audience.
-An example of a theme is the elusive nature of the American Dream, as portrayed by F.
-Scott Fitzgerald in “The Great Gatsby” (Fitzgerald and Reynolds, 1993).
-2.4. Setting
-The setting is an element that encompasses the time, place, and context in which the story unfolds (Truby, 2007).
-It includes the physical and temporal surroundings inhabited by the characters, as well as the social or cultural conventions that influence their actions and decisions.
-The setting can act dynamically, almost like a character itself, possessing specific traits and undergoing changes that directly affect the plot and contribute to the story’s development.
-This dynamic nature of the setting allows it to evolve over the course of the story, reflecting and impacting the narrative’s progression.
-The Hogwarts School of Witchcraft and Wizardry in the “Harry Potter” series (Rowling, 2015) is an example of an evolving setting that influences both plot and character development and impacts the progression of the story.
-2.5. Plot
-The plot is the sequence of events, the plot points, that propels the story from beginning to middle to end (Dibell, 1999).
-Each plot point is a moment of change that influences the understanding, decisions, and actions of the characters.
-This progression is often closely tied to the protagonists encountering or realizing a conflict, and their subsequent efforts to resolve it.
-This journey creates emotional stakes for both the characters and the audience, generating a sense of suspense.
-The plot’s movement is integral to the narrative, as it not only advances the storyline but also deepens the audience’s emotional investment in the fate of the characters.
-A single story can have multiple plots, each representing a different sequence of events or perspective.
-The plot is usually structured through cause and effect, where actions of the characters and events lead to reactions and consequences, forming a chain of scenes that collectively move the story forward.
-As an example, in “The Lord of the Rings” by J.R.R.
-Tolkien (Tolkien, 2012), the plot unfolds as a series of events beginning with Frodo inheriting the One Ring and learning about its power and dangers, followed by, among other events, the Council of Elrond and the formation of the Fellowship, Gandalf’s fall, the breaking of the Fellowship, and Frodo’s decision to, at first, continue alone.
-These plot points are influenced by character decisions and actions, which lead to various developments and challenges, enhancing the narrative.
-2.6. Mode
-The mode is a set of choices and techniques employed by the author to frame the narrative (Pinault, 1992).
-A single story can be narrated using different modes, each offering a different perspective or experience to the audience.
-The mode includes the scope of the presented information, the style and type of language used, the medium through which the story is conveyed, and the manner and extent of narrative exposition and commentary.
-A key component of the mode is the point of view or perspective from which the story is told.
-A first-person narrative is an example where a character – often the protagonist – narrates the story using personal pronouns like “I” and “me” directly engaging the audience with their perspective.
-A third-person narrative is another example that avoids personal pronouns, offering a more detached or impersonal viewpoint.
-These perspectives are examples of narration, a technique that involves the explicit telling of a story through spoken or written commentary.
-While narration is frequent in written narratives, it can be optional in other forms, like visual or interactive media.
-3. Benchmark Datasets and Evaluation Metrics
+Section 6 examines the real-world applications of VSG. Finally, Section 7 presents the key findings of this survey and proposes directions for future research.
+2 The Elements of Stories
+Stories connect events and experiences through narration, ranging from personal tales to cultural epics.
+While the terms story, narrative, and plot are often used interchangeably, they have distinct meanings: a story is the chronological sequence of events (the “what”), the narrative is how these events are presented (the “how”), and the plot is the specific sequence as experienced by the audience Simpson (2002).
+A single story can be told through multiple narratives, each offering different perspectives and experiences.
+Narratives may follow linear structures with chronological progression, or nonlinear approaches using flashbacks, flash-forwards, or parallel storylines that create complexity and require audience interpretation Lau and Chen (2008); Mitchell (2008).
+Story elements – character, conflict, theme, setting, plot, and mode – form the foundation for crafting and analyzing narratives.
+Understanding these elements is crucial for both human storytellers and automated story generation systems.
+In the following subsections, we detail each element and its significance in VSG.
+2.1 Character
+Characters are the entities through which stories unfold, with their decisions and actions propelling the plot forward DiBattista (2011).
+Through actions and speech, characters convey motives and emotions that shape audience perception and emotional connection.
+Initial character impressions influence audience expectations about future behavior.
+Characters who deviate from established patterns can create surprise or confusion, challenging audience understanding.
+Stories center around protagonists who encounter central conflicts, undergo personal growth, or acquire new knowledge.
+Antagonists add complexity and tension, often opposing the protagonist’s goals.
+Traditional narratives feature heroes battling for moral causes against villains who perpetrate evil.
+Character development and evolution throughout the narrative engages audiences and drives story progression.
+The range of character types contributes to story depth and dynamics.
+2.2 Conflict
+Conflict generates tension and drives character actions, shaping narrative trajectory and audience engagement Abbott (2008).
+How conflicts are introduced, developed, and resolved determines story progression and emotional depth.
+Conflict resolution provides closure to the central problem and marks the culmination of narrative tension.
+Conflict encompasses any tension influencing characters, from the protagonist’s main challenge to secondary internal struggles like anxiety or indecisiveness.
+These layers add complexity to character journeys and story depth.
+Longer narratives involve multiple conflicts that occur alongside the primary tension.
+Common conflict types include: Character vs.
+Character (opposition), Character vs.
+Nature (natural forces), Character vs.
+Society (social norms), Character vs.
+Unavoidable Circumstances (fate), and Character vs.
+Self (internal dilemmas).
+2.3 Theme
+Themes are the major underlying ideas of a story, often abstract and open to interpretation Griffith (2010).
+Unlike concrete story elements, themes invite discussion and allow for multiple interpretations and conclusions.
+They function as exploratory concepts that encourage audience reflection on deeper meanings.
+Themes can evolve as narratives unfold, revealing new dimensions and allowing audiences to develop personal interpretations of the story’s messages.
+Different audience members may reach varying conclusions that differ from or exceed the creator’s original intent.
+Understanding themes in narrative generation creates dynamic, thought-provoking storytelling experiences.
+2.4 Setting
+Setting encompasses the time, place, and context in which the story unfolds Truby (2007).
+It includes physical and temporal surroundings, as well as social or cultural conventions that influence character actions and decisions.
+Setting can act dynamically, possessing specific traits and undergoing changes that affect the plot.
+This allows it to evolve throughout the story, reflecting and impacting narrative progression.
+Hogwarts School in the “Harry Potter” series exemplifies an evolving setting that influences plot and character development Rowling (2015).
+2.5 Plot
+Plot is the sequence of events that propels the story from beginning to end Dibell (1999).
+Each plot point represents a moment of change that influences character understanding, decisions, and actions.
+This progression typically involves protagonists encountering conflict and attempting resolution, creating emotional stakes and suspense for both characters and audience.
+Plot advancement deepens audience investment in character fate.
+Stories can have multiple plots representing different event sequences or perspectives.
+Plot structure follows cause and effect, where character actions and events lead to reactions and consequences, forming a chain that moves the story forward.
+2.6 Mode
+Mode encompasses the choices and techniques authors use to frame narratives Pinault (1992).
+A single story can be narrated using different modes, each offering distinct audience perspectives and experiences.
+Mode includes information scope, language style, medium, and the extent of narrative exposition and commentary.
+A key component is point of view or perspective.
+First-person narratives use personal pronouns like “I” and “me,” directly engaging audiences with the character’s perspective.
+Third-person narratives avoid personal pronouns, offering more detached viewpoints.
+These perspectives exemplify narration techniques that explicitly tell stories through spoken or written commentary.
+While narration is common in written narratives, it can be optional in visual or interactive media.
+3 Benchmark Datasets and Evaluation Metrics
 This section presents and discusses benchmark datasets used for story generation and evaluation metrics that assess the performance of story generation models.
-Although not all approaches are data-driven, some of them, especially deep learning models and other machine learning approaches, rely heavily on data to guide the learning process.
-Various story datasets have been employed in the development of story generation systems by the community.
-Table 1 provides a compilation of the datasets documented in the literature for both VSG and textual story generation.
-Among these datasets, VIST (Huang et al., 2016) is prominently utilized by the research community.
+Table 1 provides a compilation of the datasets documented in the literature for VSG.
+Among these datasets, VIST Huang et al.
+(2016) is prominently utilized by the research community.
 This dataset serves as a primary benchmark, and most publications in this field report their experimental results on it.
-The Visual Writing Prompts dataset (Hong et al., 2023) is noteworthy for its focus on stories that are visually grounded on the characters and objects present in the images.
-Similarly, the Writing Prompts dataset (Fan et al., 2018) and STORIUM (Akoury et al., 2020) are important because they contain long stories derived from the imagination of the users of the platforms where they were collected.
+The VIST-Edit dataset Hsu et al.
+(2019) provides human edited versions of machine generated stories, offering insights into human preferences for story quality.
+The Visual Writing Prompts dataset Hong et al.
+(2023) is noteworthy for its focus on stories that are visually grounded on the characters and objects present in the images.
 These datasets allow for the exploration of more complex and imaginative narrative structures, serving as a resource to train models to generate fictional stories.
-Dataset No. of Stories Type Notes
-NY City (Park and Kim, 2015) 11,863 visual 11,863 blog posts / 78,467 images
-Disneyland (Park and Kim, 2015) 7,717 visual 5 7,717 blog posts / 60,545 images
-ROCStories (Mostafazadeh et al., 2016) 100,000 visual 5 sentences/story.
-LSMDC (Rohrbach et al., 2016) 118,114 visual 202 movies.
-VIST (Huang et al., 2016) 50,000 visual 210,819 unique photos
-Visual Writing Prompts (Hong et al., 2023) 12,000 visual 5 to 10 images/story, grounded on the characters.
-Pororo-SV (Kim et al., 2017) 27,328 visual
-YouCook2 (Das et al., 2013) 2,000 visual Average length of 5.26 mins.
-Writing Prompts (Fan et al., 2018) 300,000 text 28.4 words/prompt, 734.5 words/story.
-CMU movie summary corpus 42,306 visual Includes box office revenue, genre, release date, runtime, character names and actor names.
-Wikipedia’s movie plots 34,886 text Summary of the movie plot with metadata.
-Sci-fi TV plot summaries (Ammanabrolu et al., 2019) 2,276 text 80+ episodes of TV Series.
-STORIUM (Akoury et al., 2020) 5,743 text 19,278 words/story. Annotations for scene entries and cards played to guide the story.
-Table 1. Story Generation Datasets.
-The evaluation metrics, used to assess the performance of VSG models, provide quantitative measures to gauge the quality and coherence of generated narratives.
-Most of these metrics were originally developed for machine translation or summarization but have been adopted by the community for other tasks such as image captioning and story generation: instead of comparing a generated text with a reference translation, the generated text is compared with a reference caption in the case of image captioning, or with a reference story in the case of story generation.
-In the following subsections, we discuss commonly used metrics, including their strengths and limitations.
-3.1. Perplexity
-Perplexity (Jurafsky and Martin, 2023) evaluates the performance of autoregressive language models.
-It measures the model’s ability to predict a sequence and is defined as the exponential average negative log-likelihood of a sequence: given a tokenized sequence $X=(x_{1},x_{2},\ldots,x_{N})$, where $N$ is the length of the sequence, the perplexity of $X$ is given by Eq.
-1.
-$P(x_{i}|x_{1},\ldots,x_{i-1})$ denotes the probability of token $x_{i}$ given the preceding token sequence $x_{1},\ldots,x_{i-1}$ as predicted by the model.
-Perplexity can be interpreted as the geometric mean of the inverse probability of the word sequence, indicating how well the model predicts each subsequent token in the sequence.
-Lower values indicate better predictive performance, meaning that the model is more effective in capturing the underlying structure of the language.
-(1) $\text{Perplexity}(X)=\sqrt[N]{\prod_{i=1}^{N}\frac{1}{P(x_{i}|x_{1},\ldots,x_{i-1})}},$
-3.2. BLEU
-Bilingual Evaluation Understudy (BLEU) (Papineni et al., 2002) evaluates the quality of generated text by comparing it with one or more reference texts.
-It was developed for machine translation tasks, but has been adopted for various natural language generation tasks including VSG.
-BLEU outputs a score between 00 and $1$, indicating the similarity between the generated and reference texts.
-Higher scores reflect better alignment with the reference: a score of $1$ signifies perfect agreement.
-BLEU compares n-grams in the candidate text with the n-grams in the reference texts, computing the precision of these matches, which are independent of their position in the text.
-The general idea is that the greater the number of matching n-grams, the higher the quality of the text.
-To mitigate the issue of favoring shorter texts (since they can potentially have higher precision), BLEU incorporates a brevity penalty.
-This penalty balances precision and recall, ensuring that the metric does not disproportionately favor shorter stories.
-In Eq.
-2, BP is the brevity penalty (Eq.
-3), $w_{n}$ is the weight for each n-gram, $p_{n}$ is the n-gram precision, and $N$ is typically up to four, accounting for the use of unigrams, bigrams, trigrams, and four-grams.
-(2) BLEU $=\text{BP}\cdot\exp\left(\sum_{n=1}^{N}w_{n}\log p_{n}\right)$
-(3) BP =$=\begin{cases}1&\text{if }c>r\\ e^{(1-r/c)}&\text{if }c\leq r\end{cases}$
-While widely used, BLEU has limitations in evaluating stories involving word variations, synonyms, and creative text.
-Its reliance on exact n-gram matches fails to capture semantic equivalence when different wording or phrasing is used, potentially penalizing stories that are semantically accurate but stylistically different.
-This results in reduced effectiveness in assessing linguistic richness, idiomatic expressions, and the subtleties of creative language use.
-3.3. Self-BLEU
-Self-BLEU (Zhu et al., 2018) evaluates diversity within a set of generated data.
-While the original BLEU was designed to assess the similarity between two texts, Self-BLEU adapts this concept to measure how much each sentence within a generated text resembles the rest of the text.
-An individual sentence is treated as the hypothesis and all other companion sentences are considered as the reference set.
-The BLEU score is calculated for each sentence against the aggregated reference set, and the average of these scores constitutes the Self-BLEU score for the entire document.
-In Eq.
-4, $N$ is the number of sentences, $s_{i}$ is the $i$-th sentence, and $S\setminus\{s_{i}\}$ denotes the set of all sentences excluding $s_{i}$.
-[…]
-A higher Self-BLEU score indicates a lower diversity among the generated sentences, suggesting a potential mode collapse issue within the generative model.
-Mode collapse occurs when the model produces a narrow range of outputs, repeatedly generating similar types of data while failing to capture the full diversity of the target distribution.
-Conversely, a lower Self-BLEU score suggests greater diversity and a healthier, more varied generation process.
-This is particularly relevant for story generation models where diversity in narrative output is crucial for creating engaging and unique stories.
-High Self-BLEU scores might signal a lack of creativity and variability in the stories produced, whereas lower scores indicate a model’s effectiveness in generating rich, varied narratives.
-3.4. METEOR
-Metric for Evaluation of Translation with Explicit ORdering (METEOR) (Banerjee and Lavie, 2005) was designed to evaluate machine translation outputs, addressing some limitations of BLEU.
-It focuses on sentence-level evaluation, aiming to achieve higher correlation with human judgments.
-The metric relies on the harmonic mean of unigram precision and recall, with emphasis placed on recall to better align with human evaluative tendencies.
-The algorithm initiates by aligning the candidate text with the reference text, focusing on unigram mappings.
-These mappings are constrained such that each unigram in the candidate text aligns with zero or one unigram in the reference text.
-The alignment process prioritizes configurations with the fewest crossing lines to ensure clarity and coherence.
-Precision $P={m}/{w_{t}}$ and recall $R={m}/{w_{r}}$ are the ratios between $m$ (the number of matching unigrams in the candidate text) and either $w_{t}$ (the number of unigrams in the candidate text) or $w_{r}$ (the number of unigrams in the reference text).
-These metrics are integrated using the harmonic mean, emphasizing recall $F_{mean}=10\cdot P\cdot R/(R+9\cdot P)$.
-To account for the arrangement of words, METEOR introduces a penalty $p=0.5(c/u_{m})^{3}$, where $c$ is the number of chunks, and $u_{m}$ is the number of mapped unigrams.
-The final METEOR score for a candidate text is given by $M=(1-p)F_{mean}$.
-3.5. CIDEr
-Consensus-based Image Description Evaluation (CIDEr) (Vedantam et al., 2014) was developed for image captioning evaluation.
-It evaluates how well a candidate sentence $c_{i}$ aligns with a consensus among a set of reference sentences $S_{i}=\{s_{i1},\ldots,s_{im}\}$ for an image $I_{i}$.
-CIDEr involves stemming words in the sentences to their root forms.
-Each sentence is represented by its n-grams, where an n-gram $\omega_{k}$ is an ordered set of words, and n-grams from one to four words are considered.
-CIDEr evaluates the consensus by measuring how frequently n-grams in the candidate sentence appear in the reference sentences.
-It also assigns lower weights to n-grams common across all images in the dataset, as they are likely less informative.
-The Term Frequency-Inverse Document Frequency (TF-IDF) (Leskovec et al., 2014) method is employed to weight each n-gram $\omega_{k}$ based on its informativeness, as shown in Eq.
-5, where $h_{k}(s_{ij})$ is the frequency of $\omega_{k}$ in sentence $s_{ij}$, and $\Omega$ is the set of all n-grams and $|I|$ is the set of all images in the dataset.
-(5) $g_{k}(s_{ij})=\frac{h_{k}(s_{ij})}{\sum_{\omega_{l}\in\Omega}h_{l}(s_{ij})}\log\left(\frac{|I|}{\sum_{I_{p}\in I}\min(1,\sum_{q}h_{k}(s_{pq}))}\right),$
-The CIDEr score for n-grams of length $n$ is calculated using the average cosine similarity between the TF-IDF vectors of the candidate and reference sentences, capturing both precision and recall as shown in Eq.
-6, where $g_{n}(c_{i})$ and $g_{n}(s_{ij})$ are the TF-IDF vectors for n-grams of length $n$ in the candidate and reference sentences, respectively.
-$m$ is the number of reference sentences for the candidate image $i$.
-(6) $CIDEr_{n}(c_{i},S_{i})=\frac{1}{m}\sum_{j}\frac{g_{n}(c_{i})\cdot g_{n}(s_{ij})}{\|g_{n}(c_{i})\|\|g_{n}(s_{ij})\|},$
-To derive the final score, CIDEr combines scores across different n-gram lengths, as shown in Eq.
-7, where uniform weights $w_{n}=1/N$ have empirically been found to perform best, and $N$ is typically set to 4.
-(7) $CIDEr(c_{i},S_{i})=\sum_{n=1}^{N}w_{n}CIDEr_{n}(c_{i},S_{i}),$
-3.6. ROUGE
-Recall-Oriented Understudy for Gisting Evaluation (ROUGE) (Lin, 2004) is a recall-oriented measure of the overlap of n-grams between a candidate text and reference texts.
-It was developed for summarization tasks, but has been used in other natural language generation tasks.
-ROUGE-N, one of the most common variations of the metric, is the ratio of the number of n-grams co-occurring in the candidate text and the reference texts to the total number of n-grams in the reference texts, as shown in Eq.
-8, where $\text{gram}_{n}$ is the n-gram being considered, and $\text{Count}_{\text{match}}(\text{gram}_{n}(S))$ is the maximum number of n-grams co-occurring in a candidate and a set of reference texts.
-[…]
-As the number of reference texts increases, the denominator in the ROUGE-N formula also increases, which is intuitive since adding more references expands the space of acceptable texts, enhancing the robustness of the evaluation.
-For multiple references, ROUGE-N is calculated as the maximum of the pairwise ROUGE-N scores between the candidate text and each reference text (Eq.
-9, where $r_{i}$ is a reference text, and $s$ is the candidate text).
-[…]
-This approach is also extended to other ROUGE metrics such as ROUGE-L (based on the longest common subsequence), ROUGE-W (considers word sequences with gaps), and ROUGE-S (based on co-occurrence statistics of skip-bigrams, i.e., pairs of words in sentence order, allowing arbitrary gaps between them).
-3.7. BERTScore
-BERTScore (Zhang et al., 2019) is a metric that uses contextual embeddings to evaluate the similarity between a reference sentence $x=\{x_{1},\ldots,x_{k}\}$ and a candidate sentence $\hat{x}=\{\hat{x}_{1},\ldots,\hat{x}_{l}\}$.
-In story generation, $x$ and $\hat{x}$ are the reference and candidate stories, respectively.
-Each token in $x$ and $\hat{x}$ is a contextual embedding obtained from models trained on language modeling objectives, enabling the generation of distinct vectors for the same word in different contexts.
-The similarity between tokens is computed using cosine similarity.
-For pre-normalized vectors, this calculation simplifies to the inner product of the embeddings, that is, $\text{similarity}(x_{i},\hat{x}_{j})=x_{i}^{\top}\hat{x}_{j}$.
-BERTScore computes recall $R_{\text{BERT}}$, precision $P_{\text{BERT}}$, and F1 score $F_{\text{BERT}}$ by matching each token in one sentence to the most similar token in the other sentence.
-These scores are defined as shown in Eqs.
-1011, and 12, respectively.
-(10) $R_{\text{BERT}}=\frac{1}{|x|}\sum_{x_{i}\in x}\max_{\hat{x}_{j}\in\hat{x}}x_{i}^{\top}\hat{x}_{j}$
-(11) $P_{\text{BERT}}=\frac{1}{|\hat{x}|}\sum_{\hat{x}_{j}\in\hat{x}}\max_{x_{i}\in x}x_{i}^{\top}\hat{x}_{j}$
-(12) $F_{\text{BERT}}=\frac{2\cdot P_{\text{BERT}}\cdot R_{\text{BERT}}}{P_{\text{BERT}}+R_{\text{BERT}}}$
-BERTScore can incorporate importance weighting using Inverse Document Frequency (IDF) scores, emphasizing rare words which are often more indicative of sentence similarity.
-Eq.
-13 shows the recall with IDF weighting.
-(13) $R_{\text{BERT}}^{\text{idf}}=\frac{\sum_{x_{i}\in x}\text{idf}(x_{i})\max_{\hat{x}_{j}\in\hat{x}}x_{i}^{\top}\hat{x}_{j}}{\sum_{x_{i}\in x}\text{idf}(x_{i})}$
-To make BERTScore more interpretable, it is rescaled against an empirical lower bound $b$, derived from averaging BERTScores computed on randomly paired sentences from a large corpus.
-Eq.
-14 shows the rescaled BERTScore $\hat{R}_{\text{BERT}}$.
-(14) $\hat{R}_{\text{BERT}}=\frac{R_{\text{BERT}}-b}{1-b}$
-3.8. Discussion
-These metrics guide the development and refinement of visual story generation models by offering insights into their linguistic fluency, relevance to visual content, and overall narrative coherence.
+Dataset No. of Stories Notes
+NY City Park and Kim (2015) 11,863 11,863 blog posts / 78,467 images
+Disneyland Park and Kim (2015) 7,717 5 7,717 blog posts / 60,545 images
+ROCStories Mostafazadeh et al. (2016) 100,000 5 sentences/story.
+LSMDC Rohrbach et al. (2016) 118,114 202 movies.
+VIST Huang et al. (2016) 50,000 210,819 unique photos
+VIST-Edit Hsu et al. (2019) 14,905 Human-edited versions of 2,981 machine-generated stories
+Visual Writing Prompts Hong et al. (2023) 12,000 5 to 10 images/story, grounded on the characters.
+Pororo-SV Kim et al. (2017) 27,328
+YouCook2 Das et al. (2013) 2,000 Average length of 5.26 mins.
+CMU movie summary corpus 42,306 Includes box office revenue, genre, release date, runtime, character names and actor names.
+Table 1: Visual Story Generation Datasets.
+The evaluation metrics, used to assess the performance of VSG models, provide quantitative measures of the quality of generated narratives.
+Most of these metrics were originally developed for machine translation or summarization but have been adopted by the community for other tasks such as image captioning and story generation: instead of comparing a generated text with a reference translation or summary, the generated text is compared with a reference caption in the case of image captioning, or with a reference story in the case of story generation.
+These metricsa include Bilingual Evaluation Understudy (BLEU) Papineni et al.
+(2002), Self-BLEU Zhu et al.
+(2018), Metric for Evaluation of Translation with Explicit ORdering (METEOR) Banerjee and Lavie (2005), Consensus-based Image Description Evaluation (CIDEr) Vedantam et al.
+(2014), Recall-Oriented Understudy for Gisting Evaluation (ROUGE) Lin (2004), Perplexity Jurafsky and Martin (2023) and Bert Score Zhang et al.
+(2019).
+These metrics guide the development and refinement of VSG models by offering insights into their linguistic fluency, textual diversity, and mostly lexical overlap with reference texts.
 They focus on specific linguistic and content-related aspects, providing quantitative measures for precision, recall, and consensus.
-They are, however, limited by their focus on surface-level textual similarities rather than deep narrative alignment: since a story can be retold in various ways, employing different perspectives while preserving the original message, the corresponding generated text might deviate significantly from the reference text in terms of specific wording.
+However, they are limited by their focus on surface-level textual similarities rather than deep narrative alignment: since a story can be retold in various ways, employing different perspectives while preserving the original message, the corresponding generated text might deviate significantly from the reference text in terms of specific wording.
 Huang et al.
-(Huang et al., 2016) conducted an analysis to evaluate how well BLEU and METEOR correlate with human judgment in the context of VSG according to different metrics.
+Huang et al.
+(2016) conducted an analysis to evaluate how well BLEU and METEOR correlate with human judgment in the context of VSG according to different metrics.
 The results in Table 2 show that METEOR, which places a strong emphasis on paraphrasing, has higher correlation with human judgment than BLEU.
@@
 Kendall 0.14 (1.0e-33) 0.06 (8.7e-08)
-Table 2. Correlations of automatic scores against human judgements, with p-values in parentheses (Huang et al., 2016).
+Table 2: Correlations of automatic scores against human judgements, with p-values in parentheses Huang et al. (2016).
 To evaluate VSG models, additional metrics that capture the creative aspects of storytelling may be necessary.
@@
 Coherence metrics could delve deeper into the logical flow of events, ensuring that generated narratives maintain a consistent and plausible storyline.
-The proliferation of diverse LLMs, such as the GPT (OpenAI, 2024c, b, a), Llama (Touvron et al., 2023; Meta, 2024), Mistral (team, 2023), and Claude (Anthropic, 2024b, a, 2024) families, has led to the development of systematic scoring systems, aiming for objective assessments across a wide array of language understanding tasks.
-These scoring systems include Multi-turn Benchmark (MT-Bench) (Zheng et al., 2023) and Chatbot Arena (Zheng et al., 2023) briefly described below.
+The proliferation of diverse Large Language Models (LLMs), such as the GPT OpenAI (2024, 2024a, 2024b, 2024c), Llama Touvron et al.
+(2023); Meta (2024), Mistral team (2023); Mistral AI (2024), and Claude Anthropic (2025, 2024a, 2024b, 2024) families, has led to the development of systematic scoring systems, aiming for objective assessments across a wide array of language understanding tasks.
+These scoring systems include Multi-turn Benchmark (MT-Bench) Zheng et al.
+(2023) and Chatbot Arena Zheng et al.
+(2023) briefly described below.
 While they were initially developed for the evaluation of chatbots by human evaluators, they have been adapted to operate using a LLM as the evaluator achieving results close to those obtained by human evaluators.
-As LLMs continue to improve, they show potential in replacing human annotators in many tasks (Gilardi et al., 2023).
-Metrics to evaluate stories could be developed based on LLMs (Zheng et al., 2023), possibly following the underlying principles discussed above.
+As LLMs continue to improve, they show potential in replacing human annotators in many tasks Gilardi et al.
+(2023).
+Metrics to evaluate stories could be developed based on LLMs Zheng et al.
+(2023), possibly following the underlying principles discussed above.
 MT-Bench is a benchmark tailored to test the multi-turn conversation and instruction-following capabilities of LLM-based chat assistants.
@@
 In the MT-Bench evaluation process, the user is presented with two distinct conversations, each generated by a different LLM-based chat assistant and is then tasked with deciding which assistant, A, B, or indicating a tie, better followed the instructions and answered the questions.
+For VSG evaluation, this framework could be adapted by using an LLM as an automated judge to compare stories generated by different models, evaluating them based on criteria such as narrative coherence, visual grounding, and creativity.
 Chatbot Arena introduces a crowd-sourcing approach, fostering anonymous battles between LLM-based chatbot models.
 Users interact with two anonymous models simultaneously, posing identical questions to both and subsequently voting for the model that provides the preferred response.
-The Elo rating system (Zheng et al., 2023) is employed to rank the models based on their performance in these battles.
+The Elo rating system Zheng et al.
+(2023) is employed to rank the models based on their performance in these battles.
 This rating system assigns a numerical score to each model based on their battle outcomes, adjusting these scores higher or lower after each encounter to reflect their relative ability to satisfy user queries effectively compared to their opponents.
+A similar automated evaluation system for VSG could employ an LLM judge to assess pairs of generated stories given the same image sequence, ranking models based on story quality across multiple dimensions without requiring human annotators.
 In summary, while current metrics provide insights into specific dimensions of VSG, the development of new metrics may be necessary to holistically evaluate creativity, coherence, and emotional engagement, ensuring a more comprehensive understanding of the capabilities and limitations of these models.
-We anticipate that LLMs can play a role in the development of such metrics.
-4. Story Generation
-The use of Artificial Intelligence (AI) for automated story generation has been explored since the 1960s and 1970s.
-Early approaches include rule-based systems (Section 4.1), plan-based methods (Section 4.3), grammar-based systems (Section 4.2), and case-based reasoning systems (Section 4.4).
-In recent times, with the availability of large datasets and computational resources the these methods have been replaced by deep learning models: Recurrent Neural Networks (RNNs) (Section 4.5) which have been widely explored for story generation tasks; and LLMs (Section 4.6) which currently achieve state-of-the-art performance in various natural language generation tasks.
-4.1. Rule-based Methods
-Rule-based methods in narrative generation use a predefined set of rules to create stories.
-These rules govern the relationships between narrative elements, ensuring consistency within the narrative, aligning with the desired outcomes and narrative styles or genres.
-However, it is essential to acknowledge the potential limitations in flexibility and creativity, as the story’s possibilities are bounded by the predefined rules and the system’s ability to simulate complex interpersonal dynamics and character development.
-TALE-SPIN (1977) (Meehan, 1977) was an influential rule-based method that initiated its storytelling process by presenting a central problem to the main character, derived from foundational states.
-As the narrative unfolds, the character may encounter various sub-problems and unintended side effects, which also require resolution.
-The complexity of these challenges can be adjusted by the user, influencing the story’s length and intricacy.
-This capability to modulate narrative complexity allows for a range of storytelling from straightforward tales to more complex narratives, depending on the inter-character dynamics established by the user.
-A pivotal aspect of narrative construction of TALE-SPIN is its nuanced handling of character relationships which are essential for the progression and richness of the story.
-These relationships are categorized into various dimensions such as competition, dominance, familiarity, affection, trust, deceit, and indebtedness.
-These relationships not only set preconditions for the actions of the characters but also emerge as outcomes of their interactions within the narrative.
-For instance, the act of one character giving a gift to another can lead to the recipient feeling a sense of affection towards the giver, thereby influencing subsequent narrative developments.
-Fig.
-1 shows an example of a story generated by TALE-SPIN.
-[…]
-[…]
-[…]
-Figure 1. An example story by TALE-SPIN (Meehan, 1977).
-IDtension (2003) (Szilas, 2003) merges dramatic storytelling with interactive fiction.
-This architecture is designed to balance narrative consistency with the creative spontaneity essential for immersive interactive experiences.
-IDtension is composed of five foundational modules: the world of the story, narrative logic, narrative sequencer, user model, and the theater.
-Each module plays a role in crafting a narrative that responds and evolves based on user interactions, maintaining an equilibrium between narrative depth and interactive freedom.
-The world of the story module serves as the narrative’s backbone, housing the essential entities and their states, thereby setting the stage for the unfolding drama.
-The narrative logic then calculates all possible character actions derived from the narrative ecosystem.
-The narrative sequencer filters through actions to highlight those with the greatest narrative value, ensuring that each user decision is met with consequences that propel the story forward.
-The user model captures the user’s current narrative context, guiding the sequencer in tailoring the narrative flow to enhance user engagement.
-The theater is responsible for displaying the action(s), and manages the interaction between the computer and the user.
-Fig.
-2 shows an example of a story generated by IDtension.
-[…]
-[…]
-[…]
-Figure 2. An example story by IDtension (Szilas, 2003) showing the telegraphic nature of the text.
-4.2. Grammar-based Methods
-Grammar-based methods in narrative generation, a specialized subset of rule-based approaches, employ formal grammars, such as context-free grammars, to systematically construct stories.
-The vocabulary elements can be narrative elements like the ones described in the previous section.
-In contrast with rule-based methods, which emphasize logical consistency and specific narrative outcomes through predefined rules guiding character and plot interactions, these grammars consist of a set of production rules that define how various elements of a story can be combined to form coherent narratives.
-This method emphasizes linguistic correctness and structural coherence, making it suitable for applications where the narrative needs to adhere to specific stylistic or syntactic standards.
-However, the creativity and variety of the narratives generated are often limited by the constraints of the grammar used.
-Grammar-based systems are particularly effective in scenarios where the narrative needs to follow a well-defined and consistent structure.
-Different grammars have been proposed that capture narrative patterns such as the three-act structure (Aristotle et al., 1997), Propp’s morphology (Propp et al., 2010), Freytag’s pyramid (Freytag and MacEwan, 1894), the hero’s journey monomyth (Campbell, 2008), among others.
-Grammars can be designed for specific genres or domains.
-Examples of grammar-based methods include systems developed by Grimes (Ryan, 2017) and Gervás (Gervás, 2013).
-Grimes’ system was one of the pioneering approaches in the 1960s to harness computational methods for automatic story generation.
-The system would randomly select elements from Propp’s list (Propp, 1968) and then order them properly to generate a coherent story.
-This was a significant achievement, considering the computational limitations of the time.
-This system employed a grammar-based technique, drawing inspiration from Propp’s morphology, to craft structured narratives.
-Gervás (2013), decades after Grimes’ system, computationally modeled Russian folktales.
-His approach involved representing Propp’s morphology by analyzing character roles and narrative functions.
-This was implemented as a generative grammar that arranges plot elements according to a chronological sequence of narrative functions.
-4.3. Plan-based Methods
-Plan-based methods for narrative generation involve using planning algorithms to create a storyline.
-In this approach, the system formulates a narrative goal and then plans a sequence of events or actions that characters can take to achieve that goal.
-This method allows for dynamic storytelling, as the system can adapt to different narrative goals and constraints.
-However, incoherent stories can be generated if the planning algorithm is not sophisticated enough.
-Plan-based methods are useful for creating stories that require logical progression and can adapt to changing story requirements.
-Examples of plan-based methods include MEXICA (Pérez y Pérez and Sharples, 2001) and Mimesis (Young et al., 2004).
-4.3.1. MEXICA
-MEXICA (1999) (Pérez y Pérez and Sharples, 2001) used creative storytelling to generate Aztec mythology tales.
-Character-based planning drove narrative progression based on changes in character emotions extracted from a case library.
-MEXICA operates through two main processes.
-Initially, it builds knowledge from a collection of predefined stories, using this information to create schemas in its Long-Term Memory (LTM) that encapsulate content and rhetorical knowledge specific to a genre or form of writing.
-Users can add or modify stories in the system’s archive using a specialized language designed for defining these stories (PSDL) and the actions within them (SADL).
-The MEXICA storytelling core is an engagement-reflection cycle.
-In the engagement phase, MEXICA generates stories by planning a sequence of actions, each with associated post-conditions that alter the story world (e.g., emotional links between characters or changes in their physical positions).
-These post-conditions are not explicitly stated in the story but guide the narrative’s direction.
-For instance, the action “princess cures jaguar knight’s injuries” not only progresses the plot but also establishes an emotional link (gratitude) between the characters, influencing future actions.
-Characters in MEXICA possess individual “story-world contexts”, which are dynamic records of the post-conditions affecting them.
-These contexts ensure that each character’s storyline reflects their unique experiences and interactions in the narrative, allowing for complex and divergent storylines based on perspectives and relationships of the characters.
-When MEXICA reaches a narrative impasse, that is, when it cannot proceed based on the current story-world context, it enters the reflection phase.
-Here, it analyzes previous stories for inspiration to overcome the block, assessing novelty and interestingness by comparing the current story’s content and dramatic tension with those of past narratives.
-This process allows MEXICA to break impasses and continue with story generation, thereby simulating a human writer’s strategy of seeking external inspiration to resolve creative obstacles.
-Fig.
-3 shows a story generated by MEXICA.
-This example shows how limited pronoun variation and repetitive use of character names cause the narrative to be less fluid.
-[…]
-[…]
-[…]
-Figure 3. An example story by MEXICA (Pérez y Pérez and Sharples, 2001).
-4.3.2. Mimesis
-Mimesis (2004) (Young et al., 2004) addresses a fundamental challenge in game development: creating compelling and context-sensitive behaviors for characters and objects.
-Its architecture employs a structured approach to generate and execute narrative plans within interactive game environments.
-The process begins when the world planner receives a request from the game engine, detailing a narrative problem that requires resolution, such as achieving specified narrative objectives using a predetermined set of actions.
-In response, the world planner formulates a plan, a structured dataset that sequences actions for characters and other game entities to fulfill the narrative goals outlined in the request.
-Following the creation of the story world plan, the discourse planner manages non-character elements like camera movements and background music, enhancing the narrative unfolding within the game.
-The discourse planner then uses the world plan to craft a narrative plan that encapsulates all system and user activities planned in response to the game engine’s initial request.
-The final step involves the execution manager, which receives the narrative plan and builds a Directed Acyclic Graph (DAG) to map temporal dependencies among the actions.
-It then navigates through the DAG, initiating ready actions, commanding the game engine to execute them, and monitoring their outcomes to ensure successful completion or to record any failures.
-We do not include an example story of Mimesis as it is a system designed to generate stories in interactive game environments.
-4.4. Case-based Reasoning Systems
-Case-based reasoning approaches generate new stories by retrieving and adapting prior story plots stored in a memory of narrative cases.
-When a new story situation is defined, including elements like setting, characters, and author goals, this situation is matched against the narrative case library to find the most similar prior plots.
-One or more retrieved plots are adapted to the new situation by modifying characters, events and descriptions, creating a new plot that becomes the framework for the new story (it can also be added to the library for future reuse).
-This approach allows generating novel narratives by reusing and building upon successful past story structures.
-Case-based narrative systems must define how to represent, index, match, adapt, and learn from story cases.
-Large case libraries, similarity metrics, and adaptation rules enable generating a wide range of creative plot variations.
-Examples of case-based reasoning systems include MINSTREL (1993) (Turner, 1993) and Story Plot Generation (2004) (Gervás et al., 2004).
-The latter uses existing stories and case-based reasoning to create new stories based on user queries.
-MINSTREL is a creative storytelling system that generated original stories about King Arthur.
-It combines case-based goal planning with models of drama and emotional responses.
-Fig.
-4 shows an example of a story created by the system.
-[…]
-[…]
-It was the spring of 1089, and a knight named Lancelot returned to Camelot from elsewhere.
-Lancelot was hot tempered.
-Once, Lancelot lost a joust.
-Because he was hot tempered, Lancelot wanted to destroy his sword.
-Lancelot struck his sword.
-His sword was destroyed.
-One day, a lady of the court named Andrea wanted to have some berries.
-Andrea went to the woods.
-Andrea had some berries because Andrea picked some berries.
-Lancelot’s horse moved Lancelot to the woods.
-This unexpectedly caused him to be near Andrea.
-Because Lancelot was near Andrea, Lancelot saw Andrea.
-Lancelot loved Andrea.
-Some time later, Lancelot’s horse moved Lancelot to the woods unintentionally, again causing him to be near Andrea.
-Lancelot knew that Andrea kissed with a knight named Frederick because Lancelot saw that Andrea kissed with Frederick.
-Lancelot believed that Andrea loved Frederick.
-Lancelot loved Andrea.
-Because Lancelot loved Andrea, Lancelot wanted to be the love of Andrea.
-But he could not because Andrea loved Frederick.
-Lancelot hated Frederick.
-Because Lancelot was hot tempered, Lancelot wanted to kill Frederick.
-Lancelot went to Frederick.
-Lancelot fought with Frederick.
-Frederick was dead.
-Andrea went to Frederick.
-Andrea told Lancelot that Andrea was siblings with Frederick.
-Lancelot believed that Andrea was siblings with Frederick.
-Lancelot wanted to take back that he wanted to kill Frederick, but he could not because Frederick was dead.
-Lancelot hated himself.
-Lancelot became a hermit.
-Frederick was buried in the woods.
-Andrea became a nun.
-[…]
-Figure 4. An example story by MINSTREL (Turner, 1993).
-MINSTREL narratives, while coherent, lack several elements that would make them more engaging.
-The story’s phrasing is repetitive, which, along with the lack of descriptive detail, hampers the reader’s ability to fully immerse themselves in the world or connect with the characters.
-It also misses opportunities to demonstrate the emotions and motivations of the characters through their actions or dialogues, rather than through plain statements.
-4.5. Recurrent Neural Networks
-Recurrent Neural Network (RNN) are a deep learning models for sequential data processing, with applications ranging from language modeling to image captioning and text generation, that influenced later models.
-Sequential processing in RNNs involves handling sentences word by word, which, while effective for capturing short-term dependencies, poses challenges in learning long-range dependencies.
-RNNs encounter difficulties with long-range dependencies due to issues like vanishing or exploding gradients.
-This poses a limitation in capturing complex dependencies over long sequences.
-Architectures such as Long Short-Term Memory network (LSTM) (Hochreiter and Schmidhuber, 1997) and Gated Recurrent Units (GRUs) (Chung et al., 2014) were introduced to address these issues.
-Choi et al.
-(Choi et al., 2016) used RNNs to generate stories by predicting the next sentence vector based on the current input, ensuring the continuity of the narrative flow within a story.
-Jain et al.(Jain et al., 2017) used GRUs to encode sequences of variable length input descriptions and to decode them as story-like summaries.
-Fig.
-5(a) shows an example input provided to the model, and Fig.
-5(b) shows an example story generated by the model.
-RNNs base their narratives generation decisions on statistical associations derived from the datasets on which they are trained, meaning that narrative elements are learned implicitly from the training data, leading to outputs that can sometimes be unpredictable.
-In the example we can see this unpredictability in the repetition of the phrase “we also saw some lily pads .”.
-[…]
-[…]
-[…]
-(a) Input captions.
-[…]
-[…]
-[…]
-(b) Generated story.
-Figure 5. A story generated by an RNN (Jain et al., 2017).
-4.6. Transformers and Large Language Models
-Transformers (Vaswani et al., 2017) are a class of deep-learning models that represent a paradigm shift in sequential data processing and are the basis of the Large Language Models (LLMs) revolution.
-In contrast with RNNs, which process words sequentially in a recurrent way, transformers process the entire sentence simultaneously, eliminating the risk of “forgetting” past information and improving the training process efficiency.
-LLMs are composed of multiple transformer layers, which enable them to attend to information at the word, sentence, and paragraph levels.
-This multi-layered approach allows for the construction of narratives that are not only coherent but also perceived as richly detailed and creatively varied.
-LLMs also leverage their extensive training data to produce stories across a wide range of styles and genres.
-However, as with RNNs, LLMs do not explicitly learn narrative elements or the underlying structure of storytelling.
-Their ability to generate stories originates from the statistical patterns and associations found within their training corpora.
-Nevertheless, the transformer architecture’s ability to attend to information across different parts and levels of the text enables LLMs to maintain context and coherence over diverse textual structures, contributing to their effectiveness in generating narrative content.
-Bidirectional Encoder Representations from Transformers (BERT) (Devlin et al., 2019) is one of the first LLMs.
-It was widely adopted and established several state-of-the-art results for downstream tasks, even though it is now considered a small model: the base version has 110 million parameters, while current state-of-the-art models, such as Generative Pre-trained Transformer (GPT)-4 (OpenAI, 2024b) and Mixtral (team, 2023), have hundreds of billions or even trillions of parameters.
-Current LLMs, such as GPT (OpenAI, 2024b), have demonstrated proficiency in understanding context, generating coherent text, and even simulating human-like conversational abilities.
-Some of the key aspects that characterize a LLM and are crucial for their effectiveness in narrative generation are their size (number of parameters), the nature of the training data, and the context window size.
-Regarding size, larger models tend to perform better, capturing more complex language nuances.
-However, increasing model size also requires more computational resources for training and inference.
-Regarding training data, LLMs benefit from diverse and extensive training datasets: a broader range of topics and writing styles in the training data helps the model generalize well to various tasks.
-The context window size refers to the number of preceding tokens the model considers when generating the next token.
-This is a critical feature, as it determines how much of the previous text the model can “remember” and use for making predictions about what comes next.
-A larger context window allows the model to understand longer dependencies in the text, improving coherence and context awareness.
-This is especially important for tasks that involve lengthy discussions or complex narratives.
-However, larger context windows come with increased computational costs and this relationship is quadratic (Vaswani et al., 2017), posing significant challenges in terms of memory and processing power, especially for very large models and has been a key focus of recent research efforts (Ding et al., 2023).
-Several works using LLMs have been proposed for story generation.
-Tang et al.
-(Tang et al., 2022) proposed TextBox 2.0, a library that uses LLMs to generate stories.
-We do not include an example of a story generated by TextBox 2.0 as the authors did not provide one in their paper.
-Yang et al.
-(Yang et al., 2022) proposed Re3, a framework that prompts a general-purpose LLM to build a story plan consisting of a setting, characters, and a plot.
-This structured plan is generated from a short sentence describing the story, called premise.
-Story passages are then generated by repeatedly injecting contextual information from both the plan and current story state via prompt.
-This approach is one of the first approaches to explicitly model the elements character, setting, and plot in the story generation process.
-Fig.
-6 shows the initial setup and outline for an example premise.
-The resulting story incorporates all the elements from the premise, including character interactions, setting descriptions, and plot developments.
-We did not include the generated story as it occupies several pages and is available in the original paper (Yang et al., 2022).
-[…]
-[…]
-Premise: A woman discovers that her husband has been cheating on her and decides to get revenge.
-Setting: The story is set in the present day, in a small town.
-Characters:
-1. Diane Chambers is a beautiful woman in her early thirties. She has long dark hair and blue eyes. She is married to Mark Chambers and is a stay-at-home mom.
-2. Mark Chambers is a successful businessman in his early forties. He is tall and handsome, with brown hair and green eyes. He is married to Diane Chambers and is the father of two young children.
-3. Karen Johnson is a beautiful blonde women in her early thirties. She is Mark Chambers’ mistress and is also married with two young children.
-Outline:
-1. Diane Chambers discovers that her husband has been cheating on her with another woman.
-2. Diane Chambers decides to take revenge on her husband by having an affair of her own.
-[…]
-Figure 6. Re3 initial setup and outline for the shown premise (Yang et al., 2022).
-4.7. Discussion
-A critical limitation of rule-based, grammar-based, plan-based, and case-based methods is that they struggle with language variation, especially (but not only) with pronoun variation.
-This results in the frequent and repetitive use of certain structures and constituents, such as character names, leading to less fluid and less engaging narratives.
-The principle of “Show, Don’t Tell” (Dietrich and Sundell, 1983) is often underused: showing is more powerful than telling because it allows the reader to experience the story directly, rather than being told what to feel.
-However, these systems tend to rely on direct exposition to convey character emotions and story developments, rather than implying them through actions and interactions.
-This failure is not because the methods cannot solve these issues.
-Rather, they would require an impractically large number of conditions and rules to be manually encoded.
-RNNs are not immune to these challenges, but their inherent adaptability and capacity for learning from extensive datasets make them more capable of capturing the complexities of narrative storytelling.
-Likewise, LLMs are able to create fluid narratives with varied sentence structures, dialogues between characters, rich and descriptive details about the setting and characters, because they are able to learn aspects of the world from their training data.
-This allows them to generate details about the setting and characters that are not explicitly provided in the prompt.
-LLMs, however, are not without their limitations: they may suffer from hallucinations, that is, generating details that completely deviate from the original prompt, or generate contradictory or nonsensical information.
-5. From NLP to computer vision
+We anticipate that LLMs can play a role in the development of such metrics, serving as automated judges that evaluate stories based on multiple criteria.
+4 From NLP to computer vision
 This section provides an overview of areas at the intersection of computer vision and Natural Language Processing (NLP) that are relevant to the topic of VSG, namely image and video captioning and Visual Question Answering (VQA).
@@
 Some of the techniques used in these tasks have been adapted to VSG.
-Special attention is given to approaches based on deep learning architectures.
-5.1. Image and Video Captioning
+4.1 Image and Video Captioning
 The transition from image and video captioning to VSG represents a shift from describing individual images or video segments to narrating image sequences or video segments.
@@
 VSG models need to understand the temporal progression of events, causal relationships, and the ability to construct cohesive and engaging stories.
+VSG can also speculate on the motivations and emotions of characters since such interpretations do not contradict the visual content.
 Therefore, they build upon the foundational principles of image and video captioning, incorporating both the interpretation of visual data and the generation of coherent, contextually appropriate narratives.
-The tasks of image and video captioning consist in generating descriptive, human-readable sentences that accurately represent the content of an image or video, identifying the main components in a visual scene and their relationships, alongside the ability to express these details in natural language.
-Initial efforts in image and video captioning relied on template-based and rule-based approaches.
-Current deep learning models are capable of end-to-end training, resulting in a substantial enhancement in the quality of the generated captions.
-One of the most notable contributions in this domain was the application of encoder-decoder architectures, known as the ‘show and tell’ approach (Vinyals et al., 2014; Yao et al., 2015; Anderson et al., 2017).
+The tasks of image and video captioning consist in generating descriptive, human-readable sentences that represent the content of an image or video, identifying the main components in a visual scene and their relationships, alongside the ability to express these details in natural language.
+One notable contribution in this domain was the application of encoder-decoder architectures, known as the ‘show and tell’ approach Vinyals et al.
+(2014); Yao et al.
+(2015); Anderson et al.
+(2017).
 This technique employs Convolutional Neural Networks (CNNs) as encoders to extract visual features.
@@
 On the other hand, for video processing, 3D-CNNs are typically used where an additional dimension, time or depth, is taken into account to perceive the temporal dynamics in video sequences.
-Then, RNNs or their improved variants, LSTMs (Hochreiter and Schmidhuber, 1997) and GRUs (Chung et al., 2014), known for their ability to effectively model sequential data, are commonly used as decoders in image and video captioning tasks, enabling the generation of coherent and contextually meaningful textual captions and descriptions from the visual features extracted by the CNNs.
+Then, Recurrent Neural Networks (RNNs) or their improved variants, Long Short-Term Memory networks (LSTMs) Hochreiter and Schmidhuber (1997) and Gated Recurrent Units (GRUs) Chung et al.
+(2014), known for their ability to effectively model sequential data, are commonly used as decoders in image and video captioning tasks, enabling the generation of coherent and contextually meaningful textual captions and descriptions from the visual features extracted by the CNNs.
 This combination of CNNs and RNNs effectively bridges the gap between vision and language.
-In the realm of image captioning, transformers have been used as encoders to process visual features extracted from images.
-Leveraging the self-attention mechanism inherent in transformers, these models are able to capture relationships between visual elements, attending to relevant regions in the image, and highlighting visual cues and their interdependencies (Olimov et al., 2021; Radford et al., 2021; Liu et al., 2021a; Yu et al., 2019).
+In the realm of image captioning, Transformers Vaswani et al.
+(2017) have been used as encoders to process visual features extracted from images.
+Leveraging the self-attention mechanism inherent in Transformers, these models are able to capture relationships between visual elements, attending to relevant regions in the image, and highlighting visual cues and their interdependencies Olimov et al.
+(2021); Radford et al.
+(2021); Liu et al.
+(2021); Yu et al.
+(2019).
 Transformer models also demonstrate the ability to generate precise and elaborate textual descriptions, significantly improving the overall quality of image captions.
 Similarly, Transformers have made contributions to video captioning, addressing the challenges associated with understanding and describing temporal dynamics in videos.
-By treating video frames as a sequential series of images, transformers can leverage their attention mechanism to capture long-range dependencies across frames (Liu et al., 2021b; Lin et al., 2021) This enables the models to map the temporal progression of events, infer causal relationships, and grasp the underlying narrative structure within the video.
-The integration of transformers in image and video captioning has yielded improvements that surpass the performance of traditional encoder-decoder architectures (Liu et al., 2021a, b).
+By treating video frames as a sequential series of images, Transformers can leverage their attention mechanism to capture long-range dependencies across frames Liu et al.
+(2021); Lin et al.
+(2021) This enables the models to map the temporal progression of events, infer causal relationships, and grasp the underlying narrative structure within the video.
+The integration of Transformers in image and video captioning has yielded improvements that surpass the performance of traditional encoder-decoder architectures Liu et al.
+(2021, 2021).
 Transformers not only generate semantically accurate captions but also exhibit enhanced coherence and contextual relevance.
@@
 The use of Transformers in image and video captioning holds great promise, but there are challenges to be addressed.
-Efficient fusion of visual features with textual information (Radford et al., 2021), handling long-term dependencies in video sequences (Liu et al., 2021b), and ensuring temporal coherence in caption generation for longer videos are among the ongoing research topics.
-5.2. Visual Question Answering
-Visual Question Answering (VQA) combines computer vision and natural language processing to enable the extraction of information from visual content, identifying objects, scenes, actions, and comprehending context, and provide meaningful answers to questions about that content.
+Efficient fusion of visual features with textual information Radford et al.
+(2021), handling long-term dependencies in video sequences Liu et al.
+(2021), and ensuring temporal coherence in caption generation for longer videos are among the ongoing research topics.
+4.2 Visual Question Answering
+Visual Question Answering (VQA) combines computer vision and NLP to enable the extraction of information from visual content, identifying objects, scenes, actions, and comprehending context, and provide meaningful answers to questions about that content.
 Both VQA and VSG involve understanding visual content and generating meaningful responses and the developments in VQA can be directly applied in VSG tasks.
@@
 The parameters are optimized to minimize the loss between predicted answers and ground truth answers during training.
-Some examples are BUTD (Anderson et al., 2017), that uses an attention-based model that combines bottom-up and top-down mechanisms to improve performance in image captioning and visual question answering tasks, and “Show, Ask, Attend, and Answer” (Kazemi and Elqursh, 2017), that combines LSTMs with multiple attention layers.
+Some examples are BUTD Anderson et al.
+(2017), that uses an attention-based model that combines bottom-up and top-down mechanisms to improve performance in image captioning and visual question answering tasks, and “Show, Ask, Attend, and Answer” Kazemi and Elqursh (2017), that combines LSTMs with multiple attention layers.
 The transition from encoder-classifier methods to Transformer methods in VQA was driven by the Transformer’s ability to capture long-range dependencies and global contextual information.
@@
 In these VQA models, images and questions are encoded using Transformer encoders and self-attention mechanisms capture the dependencies between words and the visual features, allowing the model to focus on relevant image regions while processing the question.
-The fused encoded representations are then fed into classification layers to generate answers (Bao et al., 2022; Wang et al., 2023a, 2022, b).
-Recent advances in VQA also explore new models that approach the task as a generation problem rather than a classification one, allowing for a more flexible and diverse range of answers (Chen et al., 2022).
-6. Visual Story Generation
-The current state-of-the-art and baseline methods employed in VSG have evolved from their origins in textual story generation.
-These methods have served as foundations for the development of more sophisticated techniques and continue to provide a comparative baseline for new models and approaches.
-6.1. Visual Storytelling with Convolutional and Recurrent Neural Networks
+The fused encoded representations are then fed into classification layers to generate answers Bao et al.
+(2022); Wang et al.
+(2023, 2022, 2023).
+Recent advances in VQA also explore new models that approach the task as a generation problem rather than a classification one, allowing for a more flexible and diverse range of answers Chen et al.
+(2022).
+5 Visual Story Generation
+VSG combines computer vision and NLP to create coherent narratives from sequences of images or videos.
+This task extends beyond simple image captioning by requiring models to understand temporal relationships, infer causal connections between visual scenes, and generate engaging stories that capture the underlying narrative flow.
+Unlike video captioning, which focuses on objective description of visual content, VSG can also speculate on the motivations and emotions of characters since such interpretations do not contradict the visual content.
+The field has evolved from early encoder-decoder architectures using CNN and RNN to Transformer-based models that leverage large-scale pretraining.
+These approaches must address challenges including visual understanding, temporal coherence, and narrative creativity while maintaining grounding in the provided visual content.
+The following subsections present the key methodologies and their contributions to advancing VSG capabilities.
+5.1 Visual Storytelling with Convolutional and Recurrent Neural Networks
 Huang et al.
-(Huang et al., 2016) proposed a dataset of sequential vision-to-language mappings, SIND v.1 (later renamed to VIST), with 81,743 images in 20,211 sequences, with both descriptive (captions) and story language, and aiming at facilitating research, from understanding individual images to comprehending image sequences that narrate events over time.
-This work also proposed models for generating stories based on image sequences, using a sequence-to-sequence RNN approach, which extends the image captioning technique.
-It extracts features from each image using a pretrain CNN called AlexNET (Krizhevsky et al., 2012).
+Huang et al.
+(2016) proposed a dataset of sequential vision-to-language mappings, SIND v.1 (later renamed to VIST), with 81,743 images in 20,211 sequences, with both descriptive (captions) and story language, and aiming at facilitating research, from understanding individual images to comprehending image sequences that narrate events over time.
+This work also proposed models for generating stories based on image sequences, using a sequence-to-sequence RNN approach, which extended the image captioning technique.
+It extracts features from each image using a pretrain CNN called AlexNET Krizhevsky et al.
+(2012).
 It then encodes image sequences by using a GRU over the features of each individual image.
@@
 Four different decoding methods, shown in Table 3, were considered: beam search, greedy search, greedy search with removal of duplicates, and greedy search with removal of duplicates and incorporation of grounded words.
-Decoding Method METEOR
+Method METEOR Score
 Beam=10 23.13
-Greedy (Beam=1) 27.76
+Greedy 27.76
 Greedy - Duplicates 30.11
 Greedy - Duplicates + Grounded 31.42
-Table 3. Results of each decoding approach using evaluated using METEOR scores on the VIST (Huang et al., 2016) dataset.
+Table 3: Results of each decoding approach using evaluated using METEOR scores on the VIST Huang et al. (2016) dataset. Red indicates the lowest score, green the highest.
 Results indicate that beam search alone does not perform well in generating high-quality stories.
@@
 1 shows an example of the generated stories for an image sequence.
-Figure 7.
-Sequence of images used to generate stories using different decoding methods.
+Figure 1: Sequence of images used to generate stories using different decoding methods.
 Beam search: This is a picture of a family.
@@
 They even had a swim in the water.
-Adapted from (Huang et al., 2016).
-6.2. Visual Storytelling with Convolutional and Bidirectional Recurrent Networks Networks
+Adapted from Huang et al.
+(2016).
+5.2 Visual Storytelling with Convolutional and Bidirectional Recurrent Networks Networks
 Park et al.
-(Park and Kim, 2015) proposed Coherent Recurrent Convolutional Network (CRCN), integrating various neural network components to bridge the gap between visual content and textual narrative: VGGNet (Simonyan and Zisserman, 2015) as image encoder and Bidirectional Recurrent Networks (BRNN) to capture contextual information from both the forward and backward directions within a given text sequence.
+Park and Kim (2015) proposed Coherent Recurrent Convolutional Network (CRCN), integrating various neural network components to bridge the gap between visual content and textual narrative: VGGNet Simonyan and Zisserman (2015) as image encoder and Bidirectional Recurrent Networks (BRNN) to capture contextual information from both the forward and backward directions within a given text sequence.
 CRCN also introduces a local coherence model that focuses on maintaining coherence and context within the generated sentence sequences.
@@
 The evaluation of CRCN considered several automatic metrics and several baseline works, including one that excludes the entity coherence model.
-Table 4 shows the results for the NY City (Park and Kim, 2015) and Disneyland (Park and Kim, 2015) datasets.
+Table 4 shows the results for the NY City Park and Kim (2015) and Disneyland Park and Kim (2015) datasets.
 Both datasets were introduced by Park et al.
-(Park and Kim, 2015) and consist of image sequences with the corresponding blog posts.
+Park and Kim (2015) and consist of image sequences with the corresponding blog posts.
 In these datasets, results were superior to all baselines.
@@
 Disneyland 28.40 6.88 4.11 3.49 52.7 8.78 14.29 31.29 43.20
-Table 4.
-Performance Metrics for CRCN on New York City and Disneyland Datasets.
-Bold values represent state-of-the-art results at the time.
-B stands for BLEU, C for CIDEr, M for METEOR and R for Recall (Park and Kim, 2015).
+Table 4: Performance metrics of CRCN on New York City and Disneyland Datasets.
+Green values represent state-of-the-art results at the time.
+B stands for BLEU, C for CIDEr, M for METEOR and R for Recall Park and Kim (2015).
 Fig. 2 shows an example of a sequence of images and the corresponding generated sentences.
-Figure 8.
-Sequence of images used by CRCN a story.
+Figure 2: Sequence of images used by CRCN in a story.
 (1) One of the hallway arches inside of the library (2) As we walked through the library I noticed an exhibit called lunch hour nyc it captured my attention as I had also taken a tour of nyc food carts during my trip (3) Here is the top of the Chrysler building everyone’s favorite skyscraper in new york.
@@
 (5) We walked down fifth avenue from rockefeller centre checking out the windows in saks the designer stores and eventually making our way to the impressive new york public library.
-Adapted from (Park and Kim, 2015).
-Besides, Park et al.
-(Park and Kim, 2015) which is the first approach to use a bidirectional RNN for visual storytelling, to the best of our knowledge, XE (Wang et al., 2018) and AREL (Wang et al., 2018) have also proposed models for visual storytelling that use bidirectional RNNs on the VIST dataset.
-6.3. Transitional Adaptation of Pretrained Models for Visual Storytelling
+Adapted from Park and Kim (2015).
+In addition to the work by Park et al.
+Park and Kim (2015), which was the first to use a bidirectional RNN for visual storytelling, other works have also employed bidirectional RNNs architectures.
+XE Wang et al.
+(2018) and AREL Wang et al.
+(2018) have proposed models for visual storytelling that use bidirectional RNNs on the VIST dataset.
+5.3 Transitional Adaptation of Pretrained Models for Visual Storytelling
 Yu et al.
-(Yu et al., 2021) proposed Transitional Adaptation of Pretrained Model (TAPM), aimed at refining the generation of textual descriptions for visual content, particularly in the context of VSG tasks.
+Yu et al.
+(2021) proposed Transitional Adaptation of Pretrained Model (TAPM), aimed at refining the generation of textual descriptions for visual content, particularly in the context of VSG tasks.
 It aims at bridging the gap between pretrained language models and visual encoders.
@@
 The loss function divides sequential coherence into three components: past, current, and future matching losses.
-The past matching loss uses a Fully Connected (FC) layer $f_{p}$ to project the text representation $b_{s_{i}}$ of video $i$, drawing it nearer to the visual representation $v_{b_{i-1}}$ of the preceding video $i-1$ and distancing it from those of non-sequential videos.
+The past matching loss uses a Fully Connected (FC) layer $f_{p}$ to project the text representation $b_{s_{i}}$ of video $i$, drawing it nearer to the visual representation $v_{b_{i-1}}$ of the preceding video i1i-1 and distancing it from those of non-sequential videos.
 The future matching loss projects $b_{s_{i}}$ via a distinct FC layer $f_{f}$, aligns with the subsequent visual representation $v_{b_{i+1}}$.
@@
 The final sequential coherence loss for a given video $i$ is formulated as shown in Eq.
-15, where $\cos$ is the cosine similarity, and $j$ represents indices of incorrect matches.
-(15) $L_{i}=$ $\sum_{j\neq i-1}\max(0,1+\cos(v_{b_{j}},f_{p}(b_{s_{i}}))-\cos(v_{b_{i-1}},f_{p}(b_{s_{i}})))$
+1, where $\cos$ is the cosine similarity, and $j$ represents indices of incorrect matches.
+$L_{i}=$ $\sum_{j\neq i-1}\max(0,1+\cos(v_{b_{j}},f_{p}(b_{s_{i}}))-\cos(v_{b_{i-1}},f_{p}(b_{s_{i}})))$ (1)
 $+$ $\sum_{j\neq i}\max(0,1+\cos(v_{b_{j}},f_{c}(b_{s_{i}}))-\cos(v_{b_{i}},f_{c}(b_{s_{i}})))$
@@
 The model is then ready for the inference phase, generating captions for unseen visual inputs.
-Fig.
-9(a) shows results for TAPM against selected baselines on the LSMDC 2019 and VIST datasets.
-Fig.
-9(b) and 9(c) show human evaluation results in which TAPM surpasses adversarial baselines on LSMDC 2019.
-On VIST it surpasses the XE (Wang et al., 2018) and AREL (Wang et al., 2018) baselines in relevance, expressiveness, and concreteness.
+Table.
+6 and 6 shows results for TAPM against selected baselines on the LSMDC 2019 and VIST datasets.
+Table.
+7 show human evaluation results in which TAPM surpasses adversarial baselines on LSMDC 2019.
+On VIST it surpasses the XE Wang et al.
+(2018) and AREL Wang et al.
+(2018) baselines in relevance, expressiveness, and concreteness as shown in Table 8.
 These results highlight the strengths in word choice and contextual accuracy, showcasing its ability to capture causal relationships between images.
 However, the score is still far from human performance, indicating that there is still room for improvement.
+Models C M R
+Huang et al. Huang et al. (2016) - 31.4 -
+AREL Wang et al. (2018) 9.4 35.0 29.5
+StoryAnchor Modi and Parde (2019) 9.9 35.5 30.0
+HSRL Huang et al. (2018) 10.7 35.2 30.8
+INet Jung et al. (2020) 10.0 35.6 29.7
+TAPM 13.8 37.2 33.1
+Table 5: Quantitative results on the VIST test set. C stands for CIDEr, M for METEOR, and R for ROUGE-L. Green indicates the highest score, red the lowest.
 Models Public Test Blind Test
 C R C R
-Official Baseline (Park et al., 2018) 7.0 12.0 6.9 11.9
-XE (Wang et al., 2018) 7.2 11.5 - -
-AREL (Wang et al., 2018) 7.3 11.4 - -
+Baseline Park et al. (2018) 7.0 12.0 6.9 11.9
+XE Wang et al. (2018) 7.2 11.5 - -
+AREL Wang et al. (2018) 7.3 11.4 - -
 TAPM 10.0 12.3 8.8 12.4
-(a) Results on LSMDC 2019 public and blind test set. C stands for CIDEr and R for ROUGE.
+Table 6: Quantitative results on LSMDC 2019 public and blind test set. C stands for CIDEr and R for ROUGE. Green indicates the highest score, red the lowest. Adapted from Yu et al. (2021).
 Models Scores
 Human 1.085
-Official Baseline (Park et al., 2018) 4.015
+Official Baseline Park et al. (2018) 4.015
 TAPM 3.670
-(b) Human evaluation results on the LSMDC 2019 blind test set according to a Likert scale from 5 (worst) to 1 (best), where lower is better.
+Table 7: Human evaluation results on the LSMDC 2019 blind test set according to a Likert scale from 5 (worst) to 1 (best), where lower is better.
+Green indicates the highest score, red the lowest.
+Adapted from Yu et al.
+(2021).
 TAPM vs XE TAPM vs AREL
@@
 Concreteness 59.1 30.3 10.7 59.6 30.4 10.0
-(c) Human evaluation results on VIST. Higher is better.
-Figure 9. TAPM evaluation results. Adapted from (Yu et al., 2021).
-6.4. Combining LLMs (GPT-2) and Character Features
+Table 8: Human evaluation results on VIST. Higher is better. Green indicates the highest score, red the lowest. Adapted from Yu et al. (2021)
+5.4 Interactive and Creative Visual Storytelling
+Building upon the foundational CNN and RNN approaches, interactive and creative visual storytelling represents an important evolution in the field, introducing intermediate approaches that bridge visual comprehension with narrative creativity.
+These methods combine elements of image captioning and VQA while introducing narrative structure and creative interpretation, addressing limitations of purely technical approaches.
+Lukin et al.
+Lukin et al.
+(2018) introduced a three-module pipeline for creative visual storytelling that systematically approaches the challenge of generating narratives from visual content.
+Their pipeline consists of object identification, single-image inferencing, and multi-image narration, that serve as a preliminary design for building a creative visual storyteller.
+The approach defines computational creative visual storytelling as one with the ability to alter the telling of a story along three aspects: to speak about different environments, to produce variations based on narrative goals, and to adapt the narrative to the audience.
+This modular approach demonstrates how the foundational capabilities of image captioning and VQA can be systematically extended and combined to support more complex narrative generation tasks.
+The pipeline explicitly separates visual understanding from narrative construction, allowing for more systematic development and evaluation of each component while maintaining the overall coherence required for storytelling.
+Interactive visual storytelling has also explored user engagement and personalization in narrative generation.
+Halperin and Lukin Halperin and Lukin (2023) examined how creative visual storytelling can serve as an anthology for narrative intelligence, investigating the intersection of human creativity and automated story generation through an analysis of 100 visual stories from authors who participated in a systematic creative process of improvised story building based on image sequences.
+Their work on surreal visual storytelling Halperin and Lukin (2024) further explores how visual narrative systems can handle ambiguous or dreamlike imagery, investigating AI “hallucination” by stress-testing a visual storytelling algorithm with different visual and textual inputs designed to probe dream logic inspired by cinematic surrealism.
+These approaches highlight the importance of modularity in visual storytelling systems, where different components can be optimized independently while contributing to the overall narrative coherence.
+The integration of creative elements with systematic visual analysis provides a foundation for more sophisticated VSG systems that can balance factual visual description with imaginative narrative construction.
+5.5 Knowledge-Enhanced Visual Storytelling
+A challenge in VSG is the tendency of end-to-end approaches to produce monotonous stories with repetitive text and limited lexical diversity.
+This limitation arises because existing approaches are constrained by the vocabulary and knowledge available in single training datasets.
+To address this challenge, researchers have explored knowledge enhanced approaches that leverage external resources to enrich the story generation process.
+Hsu et al.
+Hsu et al.
+(2020) introduced KG-Story, a three-stage framework that allows VSG systems to take advantage of external knowledge graphs to produce more diverse stories.
+The framework implements a distill-enrich-generate approach: first distilling a set of representative words from input prompts, then enriching the word set using external knowledge graphs, and finally generating stories based on the enriched word set.
+This framework allows the use of external resources not only for the enrichment phase, but also for the distillation and generation phases.
+The KG-Story framework operates through three distinct stages.
+In stage 1, an image-to-term model distills representative terms from each input image, creating conceptual representations that capture the essential elements of the visual content.
+In stage 2, external knowledge graphs are utilized to identify possible connections between the extracted term sets from different images, generating enriched term paths that capture relationships and associations not explicitly visible in the images.
+In stage 3, a Transformer architecture transforms these term paths into coherent stories, incorporating techniques such as length difference positional encoding and repetition penalties to improve narrative quality.
+Evaluation results demonstrate that stories generated by KG-Story are on average ranked better than previous state-of-the-art systems in human ranking evaluations.
+The approach successfully addresses the vocabulary limitations of traditional end-to-end methods while maintaining narrative coherence.
+Interactive approaches to knowledge enhanced storytelling have also been explored.
+Hsu et al.
+Hsu et al.
+(2019) introduced Dixit, an interactive visual storytelling system that allows users to iteratively compose stories through term manipulation.
+The system extracts text terms describing objects and actions from photos, then allows users to add new terms or remove existing ones before generating stories based on these modified term sets.
+Behind the scenes, Dixit uses an LSTM model trained on image caption data to distill terms from images and utilizes a Transformer decoder to compose context coherent stories.
+This approach opens up possibilities for interpretable and controllable visual storytelling, allowing users to understand the story formation rationale and to intervene in the generation process.
+These knowledge enhanced approaches represent an important advancement in VSG, demonstrating how external knowledge resources can be systematically integrated to overcome the limitations of purely data-driven methods.
+By explicitly modeling the relationship between visual content and broader conceptual knowledge, these systems can generate more diverse, engaging, and contextually rich narratives while maintaining controllability and interpretability.
+5.6 Iterative and Planning-Based Approaches
+Traditional VSG models typically generate stories in a single forward pass.
+However, creative writers use their knowledge and worldview to put disjointed elements together to form a coherent storyline, and work and rework iteratively toward perfection.
+VSG models, however, make poor use of external knowledge and iterative generation when attempting to create stories.
+This observation has motivated the development of iterative and planning-based approaches that more closely mirror human creative writing processes.
+Hsu et al.
+Hsu et al.
+(2021) introduced PR-VIST (Plot and Rework Visual Storytelling), a framework that represents the input image sequence as a story graph in which it finds the best path to form a storyline.
+PR-VIST then takes this path and learns to generate the final story via a re-evaluating training process.
+The framework draws inspiration from the human creative writing process, which involves plotting (planning the overall narrative structure) and reworking (iteratively refining the story content).
+The PR-VIST framework operates through two main stages that correspond to the plotting and reworking phases of creative writing.
+In stage 1 (story plotting), the system constructs a story graph from the input image sequence using external knowledge graphs, including VIST and Visual Genome knowledge graphs.
+A storyline predictor model identifies the best path through this graph to form a coherent storyline, creating a structured narrative plan that connects the images through meaningful conceptual relationships.
+This plotting stage essentially creates a roadmap for the story that will be generated.
+In stage 2 (story reworking), the framework generates the actual story text based on the predicted storyline path.
+The approach uses a length controlled Transformer that is first pre-trained on the ROC Story dataset and then fine-tuned on the VIST dataset with a discriminator component to encourage higher quality story generation.
+The re-evaluating training process allows the model to iteratively improve the story generation by learning from feedback about narrative quality and coherence.
+Evaluation results demonstrate that this framework produces stories that are superior in terms of diversity, coherence, and humanness, per both automatic and human evaluations.
+An ablation study shows that both plotting and reworking contribute to the model’s superiority.
+The explicit separation of planning and generation phases allows for better control over narrative structure while maintaining the flexibility needed for creative story generation.
+This planning based approach represents an important advancement toward more human-like story generation, demonstrating how computational systems can benefit from explicitly modeling the structured creative processes that human writers naturally employ.
+5.7 Combining LLMs (GPT-2) and Character Features
 Hong et al.
-(Hong et al., 2023) proposed CharGrid (Character-Grid Transformer), a transformer model that integrates diverse input features, from global image attributes to character-specific details.
-It takes a sequence of input tokens, including global image features obtained using Swin Transformer’s (Liu et al., 2021a), character features extracted by cropping character instances, a character grid for coherence assessment, and text features introduced incrementally during processing.
+Hong et al.
+(2023) proposed CharGrid (Character-Grid Transformer), a Transformer model that integrates diverse input features, from global image attributes to character-specific details.
+It takes a sequence of input tokens, including global image features obtained using Swin Transformer’s Liu et al.
+(2021), character features extracted by cropping character instances, a character grid for coherence assessment, and text features introduced incrementally during processing.
 The trainable image and character encoders process the image feature inputs.
 The character grid is flattened and fed to a FC layer.
-The transformer module then processes these inputs, generating an output as a probability distribution over potential tokens using a pre-trained GPT-2 tokenizer.
+The Transformer module then processes these inputs, generating an output as a probability distribution over potential tokens using a pre-trained Generative Pre-trained Transformer (GPT)-2 tokenizer.
 The model architecture is shown in Fig.
-10.
-Figure 10. CharGrid architecture: the blue boxes are pre-trained components where the parameters are fixed. Adapted from (Hong et al., 2023).
+3.
+Figure 3: CharGrid architecture: the blue boxes are pre-trained components where the parameters are fixed. Adapted from Hong et al. (2023).
 During training, model parameters, excluding the vision backbone, are randomly initialized.
 The training objective involves maximizing the likelihood of image sequence-story pairs through backpropagation.
-Nucleus sampling (Holtzman et al., 2019) is employed for validation, and the METEOR score is used for evaluation.
+Nucleus sampling Holtzman et al.
+(2019) is employed for validation, and the METEOR score is used for evaluation.
 CharGrid has two additional variants: ObjGrid (Object-Grid Transformer) and EntiGrid (Entity-Grid Transformer).
@@
 CharGrid + char 47.71 25.33 11.95 5.42 33.03 25.01 4.83
-Table 5.
-Results using different input features on the test set of VWP.
+Table 9: Results using different input features on the test set of VWP.
 Global features are always included.
@@
 BLEU (B), METEOR (M), ROUGE-L (R-L), and CIDEr (C) values are averages of three runs with different random seeds.
-Adapted from (Hong et al., 2023).
-Model Grammaticality Coherence Visual Groundedness Diversity
+Green indicates the highest score, red the lowest.
+Adapted from Hong et al.
+(2023).
+Model G C VG D
 TAPM + char vs. TAPM +2.45 +1.99 +3.99* +1.69
 CharGrid vs. TAPM + char +6.49** +8.41** +6.25* +11.06**
-Table 6.
-Human binary judgments in percentage of generated stories between TAPM and TAPM with character features (TAPM + char), TAPM + char and CharGrid on the test set of VWP across four criteria: Grammaticality, Coherence, Visually Groundedness, and Diversity.
+Table 10: Human binary judgments in percentage of generated stories between TAPM and TAPM with character features (TAPM + char), TAPM + char and CharGrid on the test set of VWP across four criteria: Grammaticality (G), Coherence (C), Visual Groundedness (VG), and Diversity (D).
 The numbers are percentages.
-* means p-value $<$ 0.05.
-** means p-value $<$ 0.01.
-Adapted from (Hong et al., 2023).
-6.5. Visual Story Generation using Graphs for Event Ratiocination
+* means p-value << 0.05.
+** means p-value << 0.01.
+Adapted from Hong et al.
+(2023).
+5.8 Visual Story Generation using Graphs for Event Ratiocination
 Zheng et al.
-(Zheng et al., 2021) introduced Hypergraph-Enhanced Graph Reasoning (HEGR) for “visual event ratiocination”, a task that involves generating interpretations and narratives for events that precede, coincide with, or follow a visual scene, as well as discerning the intents of depicted characters.
+Zheng et al.
+(2021) introduced Hypergraph-Enhanced Graph Reasoning (HEGR) for “visual event ratiocination”, a task that involves generating interpretations and narratives for events that precede, coincide with, or follow a visual scene, as well as discerning the intents of depicted characters.
 The work addresses several downstream tasks, including VSG.
 The work introduces a framework to improve the integration and interpretation of visual and textual data, using hypergraphs to address challenges in multimodal interactions and temporal dynamics.
-Figure 11 shows this architecture.
-Figure 11. Architecture of HEGR by Zheng et al. Adapted from (Zheng et al., 2021).
-The method begins with feature extraction.
-Visual features are extracted using Mask R-CNN (He et al., 2017) to detect objects and their spatial characteristics within images.
-Textual features are extracted from associated descriptions or captions using BERT (Devlin et al., 2019), capturing semantic information.
+Fig.
+4 shows this architecture.
+Figure 4: Architecture of HEGR by Zheng et al. Adapted from Zheng et al. (2021).
+The method begins with feature extraction: visual features are extracted using Mask R-CNN He et al.
+(2017) to detect objects and their spatial characteristics within images.
+Textual features are extracted from associated descriptions or captions using BERT Devlin et al.
+(2019), capturing semantic information.
 In the second phase, semantic graphs are constructed for each modality: nodes represent detected features or semantic tokens, while edges depict relationships within the same modality.
@@
 This process focuses on the most relevant features for narrative generation.
-Finally, the event ratiocination generator uses the enriched feature representations to generate narratives using BART (Lewis et al., 2019).
+Finally, the event ratiocination generator uses the enriched feature representations to generate narratives using BART Lewis et al.
+(2019).
 The method was evaluated on the VIST dataset, achieving state-of-the-art performance in terms of BLEU-4, ROUGE-L and METEOR scores achieving 16.4, 37.4, and 37.8, respectively.
-6.6. Discussion
+5.9 Discussion
 The works presented in this section have made substantial contributions to the field of VSG, establishing robust benchmarks for the assessment of emerging models and methodologies.
-Specifically, the first three techniques, presented in Sections 6.16.2, and 6.3 employed a CNN encoder for visual feature extraction and a RNN decoder for language generation.
-The methods detailed in Sections 6.1 and 6.3, offer comparative analyses on the VIST dataset, with METEOR scores of 31.4 and 37.2, respectively.
+Specifically, the first three techniques, presented in Sections 5.15.2, and 5.3 employed a CNN encoder for visual feature extraction and a RNN decoder for language generation.
+The methods detailed in Sections 5.1 and 5.3, offered a comparative analyses on the VIST dataset, with METEOR scores of 31.4 and 37.2, respectively.
 Notably, the latter technique surpasses the former, which can be attributed to its transitional training approach.
@@
 Table 11 summarizes the performance of the methods described above, along with other methods from the literature that contributed to the field of VSG.
-The listed models range from earlier methods using CNNs and RNNs to more recent advances that integrate object detection models like Faster R-CNN (Ren et al., 2015) and Mask-RCNN (He et al., 2017), and LLMs like GPT-2 (Radford et al., 2019) and BART (Lewis et al., 2019).
-The first two methods, Park et al.
-(2015) (Park and Kim, 2015) and Park et al.
-(2019) (Park et al., 2018) are not comparable to the other methods as they do not use the VIST dataset.
-The remaining methods are evaluated on the VIST dataset with the exception of CharGrid (2023) (Hong et al., 2023) that is evaluated on the VWP dataset.
-The authors of CharGrid trained and evaluated the performance of TAPM on the VWP (Hong et al., 2023) dataset, making the two methods directly comparable and it is likely that CharGrid would perform slightly better than TAPM on the VIST dataset.
+The listed models range from earlier methods using CNNs and RNNs to more recent advances that integrate object detection models like Faster R-CNN Ren et al.
+(2015) and Mask-RCNN He et al.
+(2017), and LLMs like GPT-2 Radford et al.
+(2019) and BART Lewis et al.
+(2019).
+The first two methods by Park et al.
+(2015) Park and Kim (2015) and Park et al.
+(2019) Park et al.
+(2018) are not comparable to the other methods as they do not use the VIST dataset.
+The remaining methods are evaluated on the VIST dataset with the exception of CharGrid (2023) Hong et al.
+(2023) that is evaluated on the VWP dataset.
+The authors of CharGrid trained and evaluated the performance of TAPM on the VWP Hong et al.
+(2023) dataset, making the two methods directly comparable and it is likely that CharGrid would perform slightly better than TAPM on the VIST dataset.
 It is however not possible to know if CharGrid would perform better than HEGR on the VIST dataset without further experiments.
 Model Encoding Decoding Dataset B-4 R-L M
-Park et al. (2015) (Park and Kim, 2015) CNN, Bi-RNN RNN Blog Posts 3.49 - 8.78
-Park et al. (2019) (Park et al., 2018) Resnet-101/152, LSTM LSTM ActivityNet 9.91 - 16.48
-Huang et al. (2015) (Huang et al., 2016) CNN, GRU GRU VIST - - 31.4
-Xu et al. (2015) (Xu et al., 2015) VGGNet LSTM VIST - 28.94 32.98
-Yu et al. (2017) (Yu et al., 2017) ResNet-101, Bi-GRU GRU VIST - 29.53 34.12
-CST (2018) (Gonzalez-Rico and Pineda, 2018) CNN, LSTM LSTM VIST 12.7 29.2 34.4
-AREL (2018) (Wang et al., 2018) ResNet-152, Bi-GRU GRU VIST 14.1 29.5 35.0
-HSRL (2019) (Huang et al., 2018) ResNet-152 LSTM VIST 12.32 30.84 35.23
-GLACNet (2019) (Kim et al., 2018) ResNet-152, Bi-LSTM LSTM VIST - - 30.63
-StoryAnchor (2020) (Modi and Parde, 2019) ResNet-152, Bi-GRU GRU VIST 14.0 30.0 35.5
-INET (2020) (Jung et al., 2020) CNN, GRU GRU VIST 14.7 29.7 35.5
-AOG-LSTM (2023) (Liu et al., 2023) RestNet-152 LSTM VIST 12.9 30.1 36.0
-CoVS (2023) (Gu et al., 2023) RestNet-152, GRU GRU VIST 15.2 30.8 36.5
-TAPM (2021) (Yu et al., 2021) 1 Faster R-CNN, ResNet-152 GPT-2 VIST - 33.1 37.2
-HEGR (2021) (Zheng et al., 2021) 1,2 Mask-RCNN, BERT BART VIST 16.4 37.4 37.8
-TAPM (2021) (Hong et al., 2023) Faster R-CNN, ResNet-152 GPT-2 VWP 5.19 25.09 32.38
-CharGrid (2023) (Hong et al., 2023)2 Mask-RCNN, SwinTransformer GPT-2 VWP 5.42 25.01 33.03
-Table 7.
-Analysis of Visual Story Generation works.
+Park et al. (2015) Park and Kim (2015) CNN, Bi-RNN RNN Blog Posts 3.49 - 8.78
+Park et al. (2019) Park et al. (2018) Resnet-101/152, LSTM LSTM ActivityNet 9.91 - 16.48
+Huang et al. (2015) Huang et al. (2016) CNN, GRU GRU VIST - - 31.4
+Xu et al. (2015) Xu et al. (2015) VGGNet LSTM VIST - 28.94 32.98
+Yu et al. (2017) Yu et al. (2017) ResNet-101, Bi-GRU GRU VIST - 29.53 34.12
+CST (2018) Gonzalez-Rico and Pineda (2018) CNN, LSTM LSTM VIST 12.7 29.2 34.4
+AREL (2018) Wang et al. (2018) ResNet-152, Bi-GRU GRU VIST 14.1 29.5 35.0
+HSRL (2019) Huang et al. (2018) ResNet-152 LSTM VIST 12.32 30.84 35.23
+GLACNet (2019) Kim et al. (2018) ResNet-152, Bi-LSTM LSTM VIST - - 30.63
+StoryAnchor (2020) Modi and Parde (2019) ResNet-152, Bi-GRU GRU VIST 14.0 30.0 35.5
+INET (2020) Jung et al. (2020) CNN, GRU GRU VIST 14.7 29.7 35.5
+AOG-LSTM (2023) Liu et al. (2023) RestNet-152 LSTM VIST 12.9 30.1 36.0
+CoVS (2023) Gu et al. (2023) RestNet-152, GRU GRU VIST 15.2 30.8 36.5
+TAPM (2021) Yu et al. (2021) 1 Faster R-CNN, ResNet-152 GPT-2 s VIST - 33.1 37.2
+HEGR (2021) Zheng et al. (2021) 1,2 Mask-RCNN, BERT BART VIST 16.4 37.4 37.8
+TAPM (2021) Hong et al. (2023) Faster R-CNN, ResNet-152 GPT-2 VWP 5.19 25.09 32.38
+CharGrid (2023) Hong et al. (2023)2 Mask-RCNN, SwinTransformer GPT-2 VWP 5.42 25.01 33.03
+Table 11: Analysis of Visual Story Generation works.
 The columns “B-4”, “R-L”, and “M” represent the BLEU-4, ROUGE-L, and METEOR scores, respectively.
 All works, except Huang et al.
-(2015) (Huang et al., 2016) and CST (2018) (Gonzalez-Rico and Pineda, 2018), use attention mechanisms.
-All works use images as input, except ActivityNet (Fabian Caba Heilbron and Niebles, 2015) (it uses videos).
+(2015) Huang et al.
+(2016) and CST (2018) Gonzalez-Rico and Pineda (2018), use attention mechanisms.
+All works use images as input, except ActivityNet Fabian Caba Heilbron and Niebles (2015) (it uses videos).
 The mark 1 indicates works that used extra training data (direct comparison may not be fair).
-The mark 2 indicated works that explicitly model some story elements, namely the characters by the detection of objects in the case of CharGrid (2023) and HEGR (2021) and the setting in the case of HEGR (2021).
-7. Conclusions
-This survey reviewed the field of story generation, explaining its development and current methods.
-It started by explaining basic story elements and early techniques, which mostly involved manually created rules.
-We looked at the datasets and evaluation metrics used to judge the quality of the stories.
-The discussion pointed out the need for better metrics to assess complex tasks like story generation accurately.
+The mark 2 indicates works that explicitly model some story elements, namely the characters by the detection of objects in the case of CharGrid (2023) and HEGR (2021) and the setting in the case of HEGR (2021).
+6 Real-World Applications
+As VSG technologies mature, they are increasingly being deployed in real-world scenarios.
+These applications demonstrate the practical potential of VSG systems and highlight new challenges that arise when deploying these technologies in dynamic, context-rich environments.
+This section explores emerging applications that integrate visual storytelling with contextual information, location awareness, and real-time data processing.
+6.1 Location-Aware Visual Storytelling
+Suwono et al.
+Suwono et al.
+(2023) introduced location-aware visual question generation with lightweight models, representing an important advancement in spatially-grounded narrative applications.
+Their work addresses the challenge of generating contextually relevant questions and narratives that are informed by geographical location, opening new possibilities for travel documentation, educational applications, and location-based social media platforms.
+The location-aware approach integrates geographical metadata with visual content analysis to create narratives that are visually grounded and spatially contextualized.
+This integration enables the generation of stories that can reference specific landmarks, cultural contexts, and geographical features that would be meaningful to users familiar with particular locations.
+The use of lightweight models addresses practical deployment constraints, making the technology suitable for mobile applications and resource-constrained environments.
+6.2 Automotive and Real-Time Context Integration
+Belz et al.
+Belz et al.
+(2024) explored story-driven approaches in automotive contexts, investigating how real-time contextual information can enhance automated storytelling systems.
+Their research represents a step toward practical deployment of VSG in everyday scenarios, particularly in automotive systems where visual storytelling could enhance the driving experience and provide meaningful documentation of journeys.
+The automotive application domain presents challenges and opportunities for visual storytelling.
+In-car camera systems continuously capture visual data, creating an opportunity for real-time narrative generation that could help drivers and passengers understand and remember their journeys.
+The integration of real-time context information—such as speed, location, weather conditions, and traffic patterns enables the generation of richer, more informative narratives that extend beyond visual description.
+Their work demonstrates how dynamic contextual data can be incorporated into storytelling algorithms, showing measurable improvements in narrative relevance and user engagement when real-time information is provided.
+This approach could lead to applications in autonomous vehicles, where passengers might receive automatically generated travel narratives, or in fleet management systems where journey documentation becomes an automated process.
+7 Conclusions
+This survey reviewed the field of VSG, explaining its development and current methods.
+It started by explaining basic story elements, datasets and evaluation metrics used to judge the quality of the stories.
+The discussion pointed out the need for better metrics to assess complex tasks like VSG accurately.
 The document also covered related areas like image and video captioning and VQA, discussing how improvements in these fields have helped enhance VSG.
-The survey also looked at the latest developments in VSG and discussed various models that use deep learning.
-It started with basic models using simple encoder-decoder setups and went on to more advanced systems that use transformers.
-Each model was described with a focus on how they handle issues like making the story coherent, keeping it relevant to the context, and combining detailed visual and textual information.
+The discussion then moved on to the evolution of VSG methods, starting with early approaches that used CNNs and RNNs to methods based on LLMs and hypergraphs.
+Additionally, we examined emerging real-world applications that demonstrate the practical potential of these technologies in domains such as automotive systems and location-aware storytelling.
 Despite the advances, challenges remain for VSG.
@@
 We start with the ones that are more concrete and could influence the field in the short term and end with the ones that are more abstract and could potentially influence the field in the long term.
-7.1. Exploring Large Language Models
+7.1 Exploring Large Language Models
 The field of VSG using LLMs is under-explored, with only a few works delving into this domain.
-To the best of our knowledge the existing works predominantly rely on GPT-2, which is an outdated LLM when compared to models such as Mixtral (team, 2023) which is the best performing open-source model on MT-Bench or GPT-4 which is the best performing proprietary model.
+To the best of our knowledge the existing works predominantly rely on GPT-2, which is an outdated LLM when compared to models such as Qwen3-235B-A22B Yang et al.
+(2025) which is the best performing open-source model on Chatbot Arena or Gemini 2.5 Pro Comanici et al.
+(2025) which is the best performing proprietary model.
 The scarcity of the integration of LLMs with visual elements presents an opportunity for further research, exploring the potential of these models in generating more coherent and contextually relevant stories.
-7.2. Automatic Evaluation Metrics
+7.2 Automatic Evaluation Metrics
 The need for robust and efficient automatic evaluation metrics is emphasized in our survey.
@@
 Thus, while these metrics provide valuable insights into certain dimensions of text generation, they did not fully encompass the multifaceted nature of storytelling.
-As LLMs continue to improve, they show potential in replacing human annotators in many tasks (Gilardi et al., 2023) and so metrics to evaluate stories could be developed based on LLMs.
-We advance that these metrics could be based on the existing systematic scoring systems (Zheng et al., 2023) that were developed to evaluate the performance of LLM-based chat bots.
-7.3. Decomposition of Story Generation
+As LLMs continue to improve, they show potential in replacing human annotators in many tasks Gilardi et al.
+(2023) and so metrics to evaluate stories could be developed based on LLMs.
+We advance that these metrics could be based on the existing systematic scoring systems Zheng et al.
+(2023) that were developed to evaluate the performance of LLM-based chat bots.
+7.3 Decomposition of Story Generation
 Addressing the complexity of stories, we underscore the importance of decomposing the story generation process into modular components.
 Each module can focus on specific attributes such as author goals, character interactions, consistency, etc.
-This decomposition has already been explored with modeling characters in CharGrid (Yu et al., 2021) and the modeling of characters and setting in HEGR (Zheng et al., 2021).
+This decomposition has already been explored with modeling characters in CharGrid Yu et al.
+(2021) and the modeling of characters and setting in HEGR Zheng et al.
+(2021).
 We believe that explicit modeling of other story elements and abstract concepts such as emotions and culture, possibly using hypergraphs or similar structures could further enhance the quality of generated stories.
-7.4. Hybrid Systems and Knowledge Integration
+7.4 Hybrid Systems and Knowledge Integration
 Building on the decomposition concept, hybrid systems, combining planning-based techniques and machine learning approaches, present an intriguing direction.
-Hyper-graphs, as seen in HEGR (Zheng et al., 2021), could possibly be a starting point as a deep learning approach to model the planning process replacing the traditional handcrafted planning systems.
+Hyper-graphs, as seen in HEGR Zheng et al.
+(2021), could possibly be a starting point as a deep learning approach to model the planning process replacing the traditional handcrafted planning systems.
 The integration of multiple knowledge sources can enrich the diversity and quality of generated stories, paving the way for more robust and context-aware narrative creation.
+7.5 Real-World Deployment and Context Integration
+The transition from academic research to practical applications represents a crucial next step for VSG technologies.
+Recent work in location aware storytelling and automotive applications demonstrates the potential for deploying these systems in real-world scenarios.
+Future research should focus on developing robust systems that can integrate multiple contextual information sources including location, temporal data, and real-time sensor information while maintaining computational efficiency for deployment in resource constrained environments such as mobile devices and embedded systems.
+The success of these real-world applications will ultimately determine the broader impact of VSG research on society.
 Acknowledgements.
-Daniel A.
-P.
-Oliveira is supported by a scholarship granted by Fundação para a Ciência e Tecnologia (FCT), with reference 2021.06750.BD.
-Additionally, this work was supported by Portuguese national funds through FCT, with reference UIDB/50021/2020.
+[…]